### v2

# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

Abstract
Existing methods for solving Riemannian bilevel optimization (RBO) problems require prior knowledge of the problem’s first- and second-order information and curvature parameter of the Riemannian manifold to determine step sizes, which poses practical limitations when these parameters are unknown or computationally infeasible to obtain. In this paper, we introduce the Adaptive Riemannian Hypergradient Descent (AdaRHD) algorithm for solving RBO problems. To our knowledge, AdaRHD is the first method to incorporate a fully adaptive step size strategy that eliminates the need for problem-specific parameters in RBO problem resolution. We prove that AdaRHD achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity for finding an $\epsilon$-stationary point, thus matching the complexity of existing non-adaptive methods. Furthermore, we demonstrate that substituting exponential mappings with retraction mappings maintains the same complexity bound. Experiments demonstrate that AdaRHD achieves comparable performance to existing non-adaptive approaches while exhibiting greater robustness.
Key words: Adaptive algorithm, Bilevel optimization, Riemannian manifolds, Complexity
1 Introduction
Bilevel optimization has garnered significant attention owing to its diverse applications in fields such as reinforcement learning [44, 31], meta-learning [5, 60, 37], hyperparameter optimization [22, 63, 83], adversarial learning [6, 75, 76], and signal processing [45, 20]. This paper focuses on “Riemannian bilevel optimization (RBO)” [29, 19, 48], a framework that arises in applications such as Riemannian meta-learning [68], neural architecture search [65], image segmentation [24], Riemannian min-max optimization [40, 35, 88, 27, 77, 34], and Low-rank adaption [85]. The general formulation of RBO is as follows:
$\begin{array}[]{lcl}&\min\limits_{x\in{\mathcal{M}}_{x}}&F(x):=f(x,y^{*}(x)),\\ &{\rm s.t.}&y^{*}(x)=\operatorname*{arg\,min}\limits_{y\in{\mathcal{M}}_{y}}g(x,y),\end{array}$ (1)
where ${\mathcal{M}}_{x},{\mathcal{M}}_{y}$ are $d^{x}$- and $d^{y}$-dimensional complete Riemannian manifolds [48], the functions $f,g:{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\rightarrow{\mathbb{R}}$ are smooth, the lower-level function $g(x,y)$ is geodesic strongly convex w.r.t. $y$, and the upper-level function $f$ is not required to be convex.
To address Problem (1), we extend the recent advance in Euclidean adaptive bilevel optimization [81] to propose adaptive Riemannian hypergradient methods. Although certain aspects of Euclidean analysis extend naturally to Riemannian settings, the manifold’s curvature parameters introduce distortions that pose unique analytical challenges. We identify key technical difficulties in designing adaptive methods for solving RBO problems as follows:
  1. (i)
    RBO necessitates that variables are constrained on Riemannian manifolds, making it intrinsically more complex than its Euclidean counterpart. Furthermore, critical parameters such as geodesic strong convexity, Lipschitz continuity, and curvature parameters are computationally demanding to estimate over manifolds. Consequently, the geometric structure of the manifold introduces distortions that create distinct analytical challenges in the convergence analysis of adaptive algorithms.
  2. (ii)
    Compared to single-level Riemannian optimization, RBO inherently involves interdependent variable updates, resulting in complex step size selection challenges arising from coupled (hyper)gradient dynamics.
  3. (iii)
    Unlike non-adaptive RBO methods, adaptive step size strategies may exhibit divergent behavior during initial iterations, which must be rigorously addressed to enable robust theoretical convergence analysis.
In this work, we present a comprehensive convergence analysis, demonstrating our approach to resolving these challenges.
1.1 Related works
In Table 1, we summarize relevant studies related to our method, providing their applicable scenarios, adaptive or not, and the computational complexity of first- and second-order information. Note that all constants (e.g., condition numbers) are omitted from the computational complexity analysis for simplicity. For details of the related works, please refer to Appendix A.
Table 1: Comparisons of first-order and second-order complexities for reaching $\epsilon$-stationary point. Here, “Euc” and “Rie” represent the feasible region of the algorithms are Euclidean space and Riemannian manifold, respectively. The notations “Det”, “F-S”, and “Sto” represent the applicable problems as deterministic, function-sum, and stochastic, respectively. Additionally, $G_{f}$ and $G_{g}$ are the gradient complexities of $f$ and $g$, respectively. $JV_{g}$ and $HV_{g}$ are the complexities of computing the second-order cross derivative and Hessian-vector product of $g$. The notation $\tilde{\mathcal{O}}$ denotes the omission of logarithmic terms in contrast to the standard $\mathcal{O}$ notation. Furthermore, the notation “NA” indicates that the corresponding complexity is not applicable.
Methods Space Adaptive Type $G_{f}$ $G_{g}$ $JV_{g}$ $HV_{g}$
D-TFBO [81] Euc $\bm{\mathcal{\checkmark}}$ Det $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$
S-TFBO [81] $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-HINV [29] Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ NA
RHGD-CG [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-NS [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RHGD-AD [29] $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RSHGD-HINV [29] F-S $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ NA
RieBO [48] Rie Det $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
RieSBO [48] Sto $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon^{2})$ $\tilde{\mathcal{O}}(1/\epsilon^{2})$
RF2SA [19] Rie Det $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{3/2})$ NA NA
RF2SA [19] Sto $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ $\tilde{\mathcal{O}}(1/\epsilon^{7/2})$ NA NA
AdaRHD-GD (Ours) Rie $\bm{\mathcal{\checkmark}}$ Det $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$
AdaRHD-CG (Ours) $\mathcal{O}(1/\epsilon)$ $\mathcal{O}(1/\epsilon^{2})$ $\mathcal{O}(1/\epsilon)$ $\tilde{\mathcal{O}}(1/\epsilon)$
1.2 Contributions
This paper introduces the Adaptive Riemannian Hypergradient Descent (AdaRHD) algorithm, the first method to incorporate a fully adaptive step size strategy, eliminating the need for parameter-specific prior knowledge for RBO. The contributions of this work are summarized as follows:
  1. (i)
    We introduce two strategies, including Riemannian gradient descent and tangent space conjugate gradient, to compute the approximate Riemannian hypergradient and establish error bounds between the exact and approximate hypergradients;
  2. (ii)
    We develop an adaptive algorithm for solving RBO problems, eliminating reliance on prior knowledge of strong convexity, Lipschitzness, or curvature parameters. Our method achieves a convergence rate matching those of non-adaptive parameter-dependent approaches. Furthermore, by replacing exponential mappings with computationally efficient retraction mappings, our algorithm maintains comparable convergence guarantees while reducing computational costs.
  3. (iii)
    For solving smooth and strongly convex problems on Riemannian manifolds (the lower-level problem of Problem (1)), our proposed method is the first adaptive Riemannian gradient descent algorithm with a non-asymptotic convergence rate to our knowledge.
2 Preliminaries
In this section, we review standard definitions and preliminary results on Riemannian manifolds. All results presented here are available in the literature [47, 7, 29, 48], we restate them for conciseness.
We first establish essential properties and notation for Riemannian manifolds. The definition of a Riemannian manifold has been conducted in the literature [47, 7]. The Riemannian metric on a Riemannian manifold ${\mathcal{M}}$ at $x\in{\mathcal{M}}$ is denoted by $\langle\cdot,\cdot\rangle_{x}:T_{x}{\mathcal{M}}\times T_{x}{\mathcal{M}}\rightarrow{\mathbb{R}}$, with the induced norm on the tangent space $T_{x}{\mathcal{M}}$ defined as $\|u\|_{x}=\sqrt{\langle u,u\rangle_{x}}$ for an $u\in T_{x}{\mathcal{M}}$. We then recall the definitions of exponential mapping and parallel transport. For an tangent vector $u\in T_{x}{\mathcal{M}}$ and a geodesic $c:[0,1]\rightarrow{\mathcal{M}}$ satisfying $c(0)=x$, $c(1)=y$, and $c^{\prime}(0)=u$, the exponential mapping ${\rm Exp}_{x}:T_{x}{\mathcal{M}}\rightarrow{\mathcal{M}}$ is defined as ${\rm Exp}_{x}(u)=y$. The inverse of the exponential mapping, ${\rm Exp}_{x}^{-1}:{\mathcal{M}}\rightarrow T_{x}{\mathcal{M}}$, is called logarithm mapping [7, Definition 10.20], and the Riemannian distance between two points $x,y\in{\mathcal{M}}$ is given by $d(x,y)=\|{\rm Exp}_{x}^{-1}(y)\|_{x}=\|{\rm Exp}_{y}^{-1}(x)\|_{y}$. Parallel transport ${\mathcal{P}}_{x_{1}}^{x_{2}}:T_{x_{1}}{\mathcal{M}}\rightarrow T_{x_{2}}{\mathcal{M}}$ is a linear operator that preserves the inner product structure, satisfying $\langle u,v\rangle_{x_{1}}=\langle{\mathcal{P}}_{x_{1}}^{x_{2}}u,{\mathcal{P}}_{x_{1}}^{x_{2}}v\rangle_{x_{2}}$ for all $u,v\in T_{x_{1}}{\mathcal{M}}$.
We now define key properties of functions on Riemannian manifolds. The Riemannian gradient ${\mathcal{G}}f(x)\in T_{x}{\mathcal{M}}$ of a differentiable function $f\colon{\mathcal{M}}\to{\mathbb{R}}$ is the unique vector satisfying
$\langle{\mathcal{G}}f(x),u\rangle_{x}={\mathrm{D}}f(x)[u],\quad\forall u\in T_{x}{\mathcal{M}},$
where ${\mathrm{D}}f(x)[u]$ denotes the directional derivative of $f$ at $x$ along $u$ [72, 7, 48]. For twice-differentiable $f$, the Riemannian Hessian ${\mathcal{H}}f(x)$ is defined as the covariant derivative of ${\mathcal{G}}f(x)$. A function $f\colon{\mathcal{M}}\to{\mathbb{R}}$ is “geodesically (strongly) convex” if $f(c(t))$ is (strongly) convex with respect to $t\in[0,1]$ for all geodesics $c\colon[0,1]\to\Omega$, where $\Omega\subseteq{\mathcal{M}}$ is a geodesically convex set [48, Definition 4]. Specifically, if $f$ is twice-differentiable, $\mu$-geodesic strong convexity of $f$ is equivalent to ${\mathcal{H}}f(x)\succeq\mu\mathrm{Id}$, where $\mathrm{Id}$ is the identity operator. Similarly, $f$ is “geodesically $L_{f}$-Lipschitz smooth” if
$\|{\mathcal{G}}f(x)-{\mathcal{P}}_{y}^{x}{\mathcal{G}}f(y)\|_{x}\leq L_{f}d(x,y),\quad\forall x,y\in{\mathcal{M}},$
and for twice-differentiable $f$, this property is equivalent to ${\mathcal{H}}f(x)\preceq L_{f}\mathrm{Id}$.
Finally, for a bi-function $f\colon{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\to{\mathbb{R}}$, the Riemannian gradients of $f$ w.r.t. $x$ and $y$ are denoted by ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$, respectively, while the Riemannian Hessians are ${\mathcal{H}}_{x}f(x,y)$ and ${\mathcal{H}}_{y}f(x,y)$. Moreover, the Riemannian cross-derivatives ${\mathcal{G}}_{xy}^{2}f(x,y)\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ and ${\mathcal{G}}_{yx}^{2}f(x,y)\colon T_{x}{\mathcal{M}}_{x}\to T_{y}{\mathcal{M}}_{y}$, discussed in [29, 48], are linear operators. Furthermore, the operator norm of a linear operator ${\mathcal{G}}\colon T_{y}{\mathcal{M}}_{y}\to T_{x}{\mathcal{M}}_{x}$ is defined as
$\|{\mathcal{G}}\|_{x}=\sup\limits_{u\in T_{y}{\mathcal{M}}_{y},\|u\|_{y}=1}\|{\mathcal{G}}[u]\|_{x}.$
Building on the concepts of Lipschitz continuity (cf. Definition B.1) and geodesic strong convexity for functions defined on a Riemannian manifold, we present the following proposition.
Proposition 2.1 ([7, 29, 48]).
For a function $f:{\mathcal{M}}\rightarrow{\mathbb{R}}$, if its Riemannian gradient ${\mathcal{G}}f$ is $L$-Lipschitz continuous, then for all $x,y\in{\mathcal{U}}\subseteq{\mathcal{M}}$, it holds that
$f(y)\leq f(x)+\left\langle{\mathcal{G}}f(x),{\rm Exp}^{-1}_{x}(y)\right\rangle_{x}+\frac{L}{2}d^{2}(x,y).$
If $f$ is $\mu$-geodesic strongly convex, then for all $x,y\in{\mathcal{U}}$, it holds that
$f(y)\geq f(x)+\left\langle{\mathcal{G}}f(x),{\rm Exp}^{-1}_{x}(y)\right\rangle_{x}+\frac{\mu}{2}d^{2}(x,y).$ (2)
Then, for the Lipschitzness of the functions and operators, and an important result of the trigonometric distance bound over the Riemannian manifolds, we present them in Appendix B.
3 Adaptive Riemannian hypergradient descent algorithms
Standard Riemannian bilevel optimization (RBO) approaches [19, 29, 48] determine step sizes for updating the variables using problem-specific parameters such as strong convexity, Lipschitzness, and curvature constants. However, these parameters are frequently impractical to estimate or compute, posing challenges for step size determination. This limitation underscores the need for adaptive RBO algorithms that operate without prior parameter knowledge.
3.1 Approximate Riemannian hypergradient
Prior to presenting our primary algorithm, we first introduce the Riemannian hypergradient, as the core concept of our methodology relies on this construct. Following general bilevel optimization frameworks, we define the Riemannian hypergradient of $F(x)$ in Problem (1) as follows:
Proposition 3.1.
[29, 48] The Riemannian hypergradient of $F(x)$ are given by
${\mathcal{G}}F(x)={\mathcal{G}}_{x}f(x,y^{*}(x))-{\mathcal{G}}^{2}_{xy}g(x,y^{*}(x)){\mathcal{H}}_{y}^{-1}g(x,y^{*}(x))[{\mathcal{G}}_{y}f(x,y^{*}(x))]\in T_{x}{\mathcal{M}}.$ (3)
The aforementioned proposition fundamentally hinges on the implicit function theorem for Riemannian manifolds [27], which necessitates the invertibility of the Hessian of the lower-level objective function $g$ w.r.t. $y$. In this paper, this requirement is satisfied as $g$ is geodesically strongly convex w.r.t. $y$. Consequently, $y^{*}(x)$ is unique and differentiable.
However, the Riemannian hypergradient defined in (3) is computationally challenging to evaluate. First, the exact solution $y^{*}(x)$ of the lower-level objective is not explicitly available mostly, necessitating the use of an approximate solution $\hat{y}$, we employ the adaptive Riemannian gradient descent method to compute this approximation. Second, calculating the Hessian-inverse ${\mathcal{H}}_{y}^{-1}g(x,y^{*}(x))[{\mathcal{G}}_{y}f(x,y^{*}(x))]$ in (3) incurs prohibitive computational costs. To address this, given the approximation $\hat{y}$ of $y^{*}(x)$, since $g$ is geodetic strongly convex, we can approximate the Hessian-inverse-vector product ${\mathcal{H}}_{y}^{-1}g(x,\hat{y})[{\mathcal{G}}_{y}f(x,\hat{y})]$ by solving the linear system ${\mathcal{H}}_{y}g[v](x,\hat{y})={\mathcal{G}}_{y}f(x,\hat{y})$, which is originated from the following quadratic problem:
$\min_{v\in T_{\hat{y}}{\mathcal{M}}_{y}}R(x,\hat{y},v)=\frac{1}{2}\left\langle v,{\mathcal{H}}_{y}g(x,\hat{y})[v]\right\rangle_{\hat{y}}-\left\langle v,{\mathcal{G}}_{y}f(x,\hat{y})\right\rangle_{\hat{y}}.$ (4)
Denote $\hat{v}$ as the approximate solution of Problem (4). In this paper, we use adaptive gradient descent method (cf. Steps 11-15 of Algorithm 1) or tangent space conjugate gradient method [69, 7] (cf. Step 16) to obtain such a $\hat{v}$. Specifically, given an initialization $v^{0}\in T_{\hat{y}}{\mathcal{M}}_{y}$ and error tolerance $\epsilon_{v}$, the tangent space conjugate gradient algorithm for solving the linear system ${\mathcal{H}}_{y}g[v](x,\hat{y})={\mathcal{G}}_{y}f(x,\hat{y})$ is presented in Algorithm 2 of Appendix C.
Therefore, given the estimations $\hat{y}$ and $\hat{v}$, the approximate Riemannian hypergradient [19, 29, 48] is defined as follows,
$\widehat{{\mathcal{G}}}F(x,\hat{y},\hat{v})={\mathcal{G}}_{x}f(x,\hat{y})-{\mathcal{G}}^{2}_{xy}g(x,\hat{y})\hat{v}\in T_{x}{\mathcal{M}}_{x}.$ (5)
3.2 Adaptive Riemannian hypergradient descent algorithm: AdaRHD
Motivated by the recent adaptive methods [81] for solving general bilevel optimization, we employ the “inverse of cumulative gradient norm” strategy to design the adaptive step sizes [80, 78, 81], i.e., adapting the step sizes based on accumulated Riemannian (hyper)gradient norms.
Given the total iterations $T$, set the accuracies for solving the lower-level problem of Problem (1) and linear system as $\epsilon_{y}=1/T$ and $\epsilon_{v}=1/T$. Our Adaptive Riemannian Hypergradient Descent (AdaRHD) algorithm for solving Problem (1) is presented in Algorithm 1. For simplicity, when employing the gradient descent and conjugate gradient to solve the linear system, respectively, we denote AdaRHD as AdaRHD-GD and AdaRHD-CG, respectively.
Algorithm 1 Adaptive Riemannian Hypergradient Descent (AdaRHD)
1:Initial points $x_{0}\in{\mathcal{M}}_{x},y_{0}\in{\mathcal{M}}_{y}$, and $v_{0}\in T_{y_{0}}{\mathcal{M}}$, initial step sizes $a_{0}>0$, $b_{0}>0$, and $c_{0}>0$, total iterations $T$, error tolerances $\epsilon_{y}=\epsilon_{v}=\frac{1}{T}$.
2:for $t=0,1,2,...,T-1$ do
3:     Set $k=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
4:     while $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}>\epsilon_{y}$ do
5:         $b_{k+1}^{2}=b_{k}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}$,
6:         $y_{t}^{k+1}={\rm Exp}_{y_{t}^{k}}(-\frac{1}{b_{k+1}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,
7:         $k=k+1.$
8:     end while
9:     $K_{t}=k$.
10:     Set $n=0$ and $v_{t}^{0}={\mathcal{P}}_{y_{t-1}^{K_{t-1}}}^{y_{t}^{K_{t}}}v_{t-1}^{N_{t-1}}$ if $t>0$ and $v_{0}$ otherwise.
11:     while $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}>{\epsilon_{v}}$ do
12:         $c_{n+1}^{2}=c_{n}^{2}+\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}$,
13:         $v_{t}^{n+1}=v_{t}^{n}-\frac{1}{c_{n+1}}\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})$, $\triangleright$ Gradient descent
14:         $n=n+1.$
15:     end while
[…]
17:     $N_{t}=n,$
18:     $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}^{2}_{xy}g(x_{t},y_{t}^{K_{t}})[v_{t}^{N_{t}}],$
19:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,   $x_{t+1}={\rm Exp}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
20:end for
3.3 Convergence analysis
This section establishes the errors between the approximate Riemannian hypergradient (5) and the exact Riemannian hypergradient (3), the convergence result of Algorithm 1, and the corresponding computation complexity.
3.3.1 Definitions and assumptions
Similar to the definition of an $\epsilon$-stationary point in general bilevel optimization [24, 38, 39], the definition of an $\epsilon$-stationary point [19, 29, 48] in Riemannian bilevel optimization is given as follows.
Definition 3.1 ($\epsilon$-stationary point).
A point $x\in{\mathcal{M}}_{x}$ is an $\epsilon$-stationary point of Problem (1) if it satisfies $\|{\mathcal{G}}F(x)\|^{2}_{x}\leq\epsilon$.
We first concern the upper bound of curvature constant $\zeta(\tau,c)$ defined in Lemma B.1.
Assumption 3.1.
The manifolds ${\mathcal{M}}_{x}$ and ${\mathcal{M}}_{y}$ are complete Riemannian manifolds. Moreover, ${\mathcal{M}}_{y}$ is a Hadamard manifold whose sectional curvature is lower bounded by $\tau<0$. Moreover, for all iterates $y_{t}^{k}$ in the lower-level problem, the curvature constant $\zeta(\tau,d(y_{t}^{k},y^{*}(x_{t})))$ defined in Lemma B.1 is always upper bounded by a constant $\zeta$ for all $t\geq 0$ and $p\geq 0$.
Assumption 3.1 is commonly used in the convex Riemannain optimization, which ensures that the lower-level objective $g$ can be geodesically strongly convex [86, 48] and is essential for the convergence of the lower-level problem [29, 48].
We then adopt the following assumptions concerning the fundamental properties of the upper- and lower-level objectives of Problem (1).
Assumption 3.2.
Function $f(x,y)$ is continuously differentiable and $g(x,y)$ is twice continuously differentiable for all $(x,y)\in{\mathcal{U}}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$, and $g(x,y)$ is $\mu$-geodesic strongly convex in $y\in{\mathcal{U}}_{y}$ for any $x\in{\mathcal{U}}_{x}$.
This assumption is prevalent in (Riemannian) bilevel optimization [24, 38, 39, 51, 32, 46, 19, 29, 48]. Under this assumption, the optimal solution $y^{*}(x)$ of the lower-level problem is unique for all $x\in{\mathcal{M}}_{x}$ and differentiable [29, 48], ensuring the existence of the Riemannian hypergradient (3.1).
Assumption 3.3.
Function $f(x,y)$ is $l_{f}$-Lipschitz continuous in ${\mathcal{U}}\in{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$. The Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous in ${\mathcal{U}}$. The Riemannian gradients ${\mathcal{G}}_{x}g(x,y)$ and ${\mathcal{G}}_{y}g(x,y)$ are $L_{g}$-Lipschitz continuous in ${\mathcal{U}}$. Further, the Riemannian Hessian ${\mathcal{H}}_{y}g(x,y)$, cross derivatives ${\mathcal{G}}^{2}_{xy}g(x,y)$, ${\mathcal{G}}^{2}_{yx}g(x,y)$ are $\rho$-Lipschitz in ${\mathcal{U}}$.
Assumption 3.3 is a standard condition in the existing literature on (Riemannian) bilevel optimization [24, 38, 39, 29, 48]. These conditions ensure the smoothness of the function $F$ and the Riemannian hypergradient ${\mathcal{G}}F$.
Assumption 3.4.
The minimum of $F$ over ${\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$, denote as $F^{*}$, is lower-bounded.
Assumption 3.4 concerns the existence of the minimum of the upper-level problem, which is a minor requirement.
3.3.2 Convergence results
Denote $\hat{v}_{t}^{*}(x,y):=\operatorname*{arg\,min}_{v\in T_{y}{\mathcal{M}}_{y}}R(x,y,v)$. We can then bound the estimation error of the proposed schemes of approximated hypergradient as follows.
Lemma 3.1 (Hypergradient approximation error bound).
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. The error for the approximated Riemannian hypergradient generated by Algorithm 1 satisfies,
$\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})-{\mathcal{G}}F(x_{t})\|_{x_{t}}\leq L_{1}d\left(y^{*}(x_{t}),y_{t}^{K_{t}}\right)+L_{g}\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}},$
where $L_{1}:=L_{f}+\frac{l_{f}\rho}{\mu}+L_{g}(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})$.
Lemma 3.1 establishes that the error between the approximate and exact Riemannian hypergradients is bounded by the errors arising from the resolutions of the lower-level problem in (1) and the linear system (4).
Similar to Proposition 2 in [81], we derive upper bounds on the total iteration count required to solve both the lower-level problem in Problem (1) and the linear system (4). However, owing to the geometric structures inherent in Riemannian manifolds, these bounds cannot be directly inferred from [81, Proposition 2] and require a meticulous analysis.
Proposition 3.2.
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. Then, for any $0\leq t\leq T$, the iterations $K_{t}$ and $N_{t}$ required in Algorithm 1 satisfy:
$K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}+\frac{1}{({\mu}/{\bar{b}}-\zeta{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{\tilde{b}}{\epsilon_{y}}\right).$
AdaRHD-GD:
$N_{t}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}+\frac{\bar{c}}{\mu}\log\left(\frac{\tilde{c}}{\epsilon_{v}}\right),$
where $C_{b},C_{c}$, $\bar{b}$, $\bar{c}$, $\tilde{b}$, and $\tilde{c}$ are constants defined in Appendix G.
AdaRHD-CG:
$N_{t}\leq\frac{1}{2}{\log\left(\frac{4L_{g}^{3}l_{f}^{2}}{\mu^{3}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{L_{g}/\mu}+1}{\sqrt{L_{g}/\mu}-1}\right)}.$
Proposition 3.2 provides upper bounds for the total iterations of the resolutions of the lower-level problem in Problem (1) and the linear system 4. If $\epsilon_{y}>0$ is sufficiently small, it holds that $K_{t}=\mathcal{O}(1/\epsilon_{y})$, which matches the standard AdaGrad-Norm for solving strongly convex problems [80]. Additionally, if $\epsilon_{v}>0$ is small enough, we also have $N_{t}=\mathcal{O}(1/\epsilon_{v})$ and $N_{t}=\mathcal{O}(\log(1/\epsilon_{v}))$ when employing gradient descent and conjugate gradient methods, respectively.
Based on Proposition 3.2, we have the following convergence result of Algorithm 1.
Theorem 3.1.
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. The sequence $\{x_{t}\}_{t=0}^{T}$ generated by Algorithm 1 satisfies,
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|{\mathcal{G}}F(x_{t})\right\|^{2}_{x_{t}}\leq\frac{C}{T}=\mathcal{O}\left(\frac{1}{T}\right),$
where $C$ is a constant defined in Appendix G.
Theorem 3.1 establishes that the proposed adaptive algorithm achieves a convergence rate matching standard non-adaptive (Riemannian) bilevel algorithms (e.g., [24, 38, 39, 29, 48]), demonstrating its computational efficiency. A complexity bound for Algorithm 1 is subsequently derived.
Corollary 3.1.
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. Algorithm 1 needs $T=\mathcal{O}(1/\epsilon)$ iterations to achieve an $\epsilon$-accurate stationary point of Problem (1). The gradient complexities of $f$ and $g$ are $G_{f}=\mathcal{O}(1/\epsilon)$ and $G_{g}=\mathcal{O}(1/\epsilon^{2})$, respectively. The complexities of computing the second-order cross derivative and Hessian-vector product of $g$ are $JV_{g}=\mathcal{O}(1/\epsilon)$, $HV_{g}=\mathcal{O}(1/\epsilon^{2})$ for AdaRHD-GD, and $HV_{g}=\tilde{\mathcal{O}}(1/\epsilon)$ for AdaRHD-CG, respectively.
Corollary 3.1 demonstrates that the gradient complexity ($G_{g}$) and Hessian-vector product complexity ($HV_{g}$) of AdaRHD-GD surpass those of well-tuned Riemannian bilevel optimization methods [29, 48] by a factor of $1/\epsilon$. This gap originates from the additional iterations needed to guarantee the solution of the lower-level problem in Problem (1) and linear system 4, which is necessitated by the lack of prior knowledge regarding the strong convexity, Lipschitzness, and curvature constants.
3.4 Extend to retraction mapping
Given that retraction mapping is often preferred in practice for its computational efficiency over exponential mapping, this section demonstrates that its incorporation into Algorithm 1 achieves comparable theoretical convergence guarantees. A modified version of Algorithm 1, incorporating retraction mapping, is presented in Algorithm 3 (cf. Appendix D). To formalize this extension, the second condition of Assumption 3.1 must be modified as follows.
Assumption 3.5.
All the iterates of the lower-level problem generated by Algorithm 3 lie in a bounded set belonging to ${\mathcal{M}}_{y}$ that contains the optimal solution, i.e., for all $t\geq 0$ there exists a constant $\bar{D}>0$, such that $d(y_{t}^{k},y^{*}(x_{t}))\leq\bar{D}$ for all $k\geq 0$.
Under Assumption 3.5, the curvature constant $\zeta(\tau,d(y_{t}^{k},y^{*}(x_{t})))$ in Assumption 3.1 is bounded by its definition, thereby ensuring consistency with Assumption 3.1. Additionally, it is necessary to consider the error between the exponential and retraction mappings.
Assumption 3.6.
Given $z_{1}\in{\mathcal{U}}\in{\mathcal{M}}$ (here ${\mathcal{M}}$ can be ${\mathcal{M}}_{x}$ or ${\mathcal{M}}_{y}$) and $u\in T_{x}{\mathcal{M}}$, let $z_{2}={\rm Retr}_{z_{1}}(u)$. There exist constants $c_{u}\geq 1,c_{R}\geq 0$ such that $d^{2}(z_{1},z_{2})\leq c_{u}\|u\|_{z_{1}}^{2}$ and $\|{\rm Exp}^{-1}_{z_{1}}(z_{2})-u\|_{z_{1}}\leq c_{R}\|u\|_{z_{1}}^{2}$.
Assumption 3.6 is standard (e.g., [29, Assumption 2] and [41, Assumption 1]) in bounding the error between the exponential mapping and retraction mapping, given that retraction mapping is a first-order approximation to the exponential mapping [29].
Then, similar to Theorem 3.1, we have the following convergence result of Algorithm 3.
Theorem 3.2.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, and 3.6 hold. The sequence $\{x_{t}\}_{t=0}^{T}$ generated by Algorithm 3 staisifes,
$\frac{1}{T}\sum_{t=0}^{T-1}\left\|{\mathcal{G}}F(x_{t})\right\|^{2}_{x_{t}}\leq\frac{C_{{\rm retr}}}{T}=\mathcal{O}\left(\frac{1}{T}\right),$
where $C_{{\rm retr}}$ is a constant defined in Appendix H.
Furthermore, the complexity analysis of Algorithm 3 closely aligns with that of Algorithm 1 and is, therefore, omitted.
4 Experimental results
In this section, adhering to the experimental framework established by [29], we conduct comprehensive experiments to evaluate our algorithm against the RHGD method. Since the RieBO algorithm proposed by [48] shares the same algorithmic framework as RHGD, we categorize them as a single method for comparison. We designate RHGD-50 and RHGD-20 as configurations with maximum iteration limits for solving the lower-level problem in RHGD set to 50 and 20, respectively. For consistency, we exclusively employ Algorithm 3 in this section due to its computational efficiency in practice. Detailed experimental settings and additional experiments are provided in Appendix I.
4.1 Simple problem
In the first experiment, following [29], we consider a simple problem, which aims to determine the maximum similarity between two matrices ${\mathbf{X}}\in{\mathbb{R}}^{n\times d}$ and ${\mathbf{Y}}\in{\mathbb{R}}^{n\times r}$, where $n\geq d\geq r$, formulated as:
$\begin{array}[]{lcl}&\max\limits_{{\mathbf{W}}\in{\rm St}(d,r)}&{\rm trace}(\mathbf{M}^{*}({\mathbf{W}}){\mathbf{X}}^{\top}{\mathbf{Y}}{\mathbf{W}}^{\top})\\ &{\rm s.t.}&\mathbf{M}^{*}({\mathbf{W}})=\operatorname*{arg\,min}\limits_{\mathbf{M}\in{\mathbb{S}}_{++}^{d}}\ \ \langle\mathbf{M},{\mathbf{X}}^{\top}{\mathbf{X}}\rangle+\langle\mathbf{M}^{-1},{\mathbf{W}}{\mathbf{Y}}^{\top}{\mathbf{Y}}{\mathbf{W}}^{\top}+\lambda{\mathbf{I}}\rangle,\end{array}$
where ${\rm St}(d,r)=\{{\mathbf{W}}\in{\mathbb{R}}^{d\times r}:{\mathbf{W}}^{\top}{\mathbf{W}}={\mathbf{I}}_{r}\}$ and ${\mathbb{S}}_{++}^{d}=\{\mathbf{M}\in{\mathbb{R}}^{d\times d}:\mathbf{M}\succ 0\}$ represent the Stiefel manifold and the symmetric positive definite (SPD) matrices, respectively, and $\lambda>0$ is the regular parameter. The matrix ${\mathbf{W}}\in{\rm St}(d,r)$ aligns ${\mathbf{X}}$ and ${\mathbf{Y}}$ in a shared dimensional space, while the lower-level problem learns an appropriate geometric metric $\mathbf{M}\in{\mathbb{S}}_{++}^{d}$ [84]. Additionally, the geodesic strong convexity of the lower-level problem and the Hessian inverse expression can be found in Appendix H of [29]. In this experiment, we generate random data matrices ${\mathbf{X}}$ and ${\mathbf{Y}}$ with two sample sizes: $n=100$ and $n=1000$, where $d=50$ and $r=20$.
Figure 1: Performances of methods in $n=100$.
Figure 2: Performances of methods in $n=1000$.
Figures 2 and 2 show the evolution of the upper-level objective function (Upper Objective) over time and the associated hypergradient estimation errors (Hypergrad error) across outer iterations. Figure 2 corresponds to $n=100$, while Figure 2 represents the $n=1000$ case. These results demonstrate that our algorithm exhibits faster convergence and superior performance compared to RHGD while maintaining scalability with increasing sample dimensionality, confirming its efficiency and robustness. Moreover, although GD achieves quicker initial convergence, CG demonstrates greater robustness, as evidenced by its lower hypergradient estimation errors.
4.2 Robustness analysis
In the second experiment, also from [29, Section 4.2], we address the deep hyper-representation problem for classification, which is a subclass of hyper-representation problems [22, 63, 83, 53, 64]. In contrast to the 2-layer SPD network employed for ETH-80 dataset classification in [29], to demonstrate the efficacy of our algorithm, here we utilize a 3-layer SPD network [36] as the upper-level architecture to optimize input embeddings over the larger AFEW dataset [16], comprising seven emotion classes. The training set contains 1,747 matrices with imbalanced class distribution (267, 235, 173, 292, 342, 288, and 150 samples per class, respectively). This optimization problem is formulated as follows:
$\begin{array}[]{lcl}&\min\limits_{\mathbf{A}_{1},~{}\mathbf{A}_{2},~{}\mathbf{A}_{3}}&-\sum\limits_{i\in{\mathcal{D}}_{\rm val}}\frac{{\mathbf{y}}_{i}^{\top}\log({\rm SPDnet}(\mathbf{D}_{i};\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})\beta^{*}(\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3}))}{|{\mathcal{D}}_{\rm val}|},\\ &{\rm s.t.}&\beta^{*}(\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})=\operatorname*{arg\,min}\limits_{\beta\in{\mathbb{R}}^{{r(r+1)/2}}}-\sum\limits_{i\in{\mathcal{D}}_{\rm tr}}\frac{{\mathbf{y}}_{i}^{\top}\log({\rm SPDnet}(\mathbf{D}_{i};\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})\beta)}{|{\mathcal{D}}_{\rm tr}|}+\frac{\lambda}{2}\|\beta\|^{2},\\ &&\mathbf{A}_{1}\in{\rm St}(d,d_{1}),~{}\mathbf{A}_{2}\in{\rm St}(d_{1},d_{2}),~{}\mathbf{A}_{3}\in{\rm St}(d_{2},r),\end{array}$
where $\text{SPDnet}(\cdot;\mathbf{A}_{1},\mathbf{A}_{2},\mathbf{A}_{3})$ denotes the 3-layer SPD network with layer parameters $\mathbf{A}_{1}$, $\mathbf{A}_{2}$, and $\mathbf{A}_{3}$ respectively [36], the term ${\mathbf{y}}_{i}$ represents the one-hot encoded label., and ${\mathcal{D}}=\{\mathbf{D}_{i}\}_{i=1}^{n}$ denote a set of SPD matrices where $\mathbf{D}_{i}\in{\mathbb{S}}_{++}^{d}$. Each matrix $\mathbf{D}_{i}$ has dimensions of $400\times 400$, and we set $d_{1}=100$, $d_{2}=20$, and $r=5$.
In this study, we perform a series of experiments with varying initial step sizes to evaluate the robustness and advantages of our proposed AdaRHD algorithm. To address computational constraints, we utilize a $5\%$ subset of the AFEW dataset [16] rather than the full dataset, ensuring tractable training durations. For AdaRHD, we initialize hyperparameters $a_{0}$, $b_{0}$, and $c_{0}$ to equal values and test performance across the range $\{0.2,1,2,10,20\}$. For RHGD, we fix $\eta_{x}$ and $\eta_{y}$ to equal values and evaluate the set $\{5,1,0.5,0.1,0.05\}$. To mitigate sampling bias and validate robustness, each algorithm is executed five times with distinct random seeds, each iteration employing a unique $5\%$ data subset. This methodology enables systematic assessment of optimization sensitivity to step sizes and establishes the generalizability of our algorithm under practical constraints.
Figure 3: Epoch vs. Validation Accuracy under different initial step sizes for each algorithm. In the figures for “ARHD-X”, the labels indicate the values of $1/a_{0}=1/b_{0}=1/c_{0}$; in the figures for “RHGD-X”, the labels represent the values of $\eta_{x}=\eta_{y}$.
Table 2: Time to Reach a Specific Validation Accuracy Under Different Initial Step Sizes for Each Algorithm. The values outside the parentheses indicate the mean over five random trials, while the values inside the parentheses represent the standard deviation. “Step Size (A/R)” denotes the initial step sizes used in each method: $1/a_{0}=1/b_{0}=1/c_{0}$ in the “ARHD-X” algorithm; $\eta_{x}=\eta_{y}$ in the “RHGD-X” algorithm. “X%” indicates the time required to reach the corresponding validation accuracy.
AdaRHD-CG AdaRHD-GD RHGD-50 RHGD-20
Step Size 50% 70% 85% 50% 70% 85% 50% 70% 85% 50% 70% 85%
5 280.91 468.03 1238.98 449.27 677.38 1346.98 / / / / / /
(91.02) (127.79) (669.09) (172.40) (142.76) (287.18) / / / / / /
1 128.77 187.86 241.78 213.68 320.82 425.00 / / / / / /
(19.52) (33.53) (71.50) (31.74) (55.81) (93.22) / / / / / /
0.5 115.37 167.17 237.84 192.50 293.99 466.93 / / / / / /
(18.23) (26.20) (29.97) (39.14) (59.67) (108.76) / / / / / /
0.1 222.24 319.61 436.05 306.54 428.54 541.38 165.28 358.56 789.99 33.52 158.94 527.74
(35.19) (35.02) (51.80) (55.28) (45.77) (82.30) (76.28) (70.23) (161.80) (11.87) (114.19) (62.51)
0.05 254.24 419.30 642.35 351.56 506.29 650.03 135.89 693.97 1321.19 39.41 168.10 1036.95
(41.66) (53.76) (108.29) (35.66) (25.78) (51.24) (135.61) (88.98) (175.85) (6.92) (145.02) (131.94)
The results are presented in Figure 3 and Table 2. It can be observed that the RHGD-50 and RHGD-20 fail when the initial step sizes are set to 5, 1, or 0.5, whereas the corresponding configurations of AdaRHD remain relatively stable. Notably, even when using a step size greater than 1 (i.e., $a_{0}=b_{0}=c_{0}=0.2$), although the performance is less stable compared to other settings, AdaRHD is still able to converge effectively. Table 2 also shows that AdaRHD-CG consistently achieves 85% validation accuracy in the shortest amount of time. On the other hand, step sizes of $a_{0}=b_{0}=c_{0}=1$ or $2$ yield the best performance among all initializations for the AdaRHD variants, while the corresponding RHGD configurations fail to converge. These results further demonstrate the robustness of the proposed AdaRHD method, which significantly reduces the sensitivity to the choice of initial step size and greatly improves training efficiency. Moreover, although RHGD-20 requires the shortest time to reach 50% and 70% validation accuracies, it demands more time than our method to achieve 85% validation accuracy, demonstrating the efficiency of our approach. Meanwhile, the lower standard deviation exhibited by our method further highlights the robustness of our algorithm.
5 Conclusion
This paper proposes an adaptive algorithm for solving Riemannian bilevel optimization (RBO) problems, employing two strategies: gradient descent and conjugate gradient, to compute approximate Riemannian hypergradients. To our knowledge, this is the first fully adaptive RBO algorithm, incorporating a step size mechanism eliminating prior knowledge requirements of problem parameters. We establish that the method achieves an $\mathcal{O}(1/\epsilon)$ iteration complexity to reach an $\epsilon$-stationary point, matching the complexity of the standard non-adaptive algorithms. Additionally, we show that substituting the exponential mapping with a computationally efficient retraction mapping maintains this complexity guarantee.
Appendix
Appendix A Related works
Bilevel optimization: Bilevel optimization is rooted in the Stackelberg game [10, 6, 75, 76]. When the lower-level objective is strongly convex, various methods have been proposed [24, 12, 38, 39, 15, 51, 32, 46]. Hypergradient computation approaches include approximate implicit differentiation (AID) [17, 58]; iterative differentiation (ITD) [54, 21, 63, 25]; Neumann series (NS) [24]; and conjugate gradient (CG) [38]. Recent studies address bilevel problems with constraints on either the upper- or lower-level objectives. For example, [13, 32] focus on scenarios where constraints are imposed solely on the upper-level objective, whereas [71, 79, 82] develop methods for problems with lower-level constraints. For further reading, readers are referred to [38, 39, 46] and references therein.
Riemannian manifold optimization: Riemannian manifold optimization has attracted considerable attention due to its broad applications, including low-rank matrix completion [8, 73]; phase retrieval [4, 67]; dictionary learning [14, 66]; dimensionality reduction [30, 70, 56]; and manifold regression [49, 50]. This methodology reformulates manifold-constrained problems into unconstrained ones by treating the manifold as the ambient space and employing retractions to compensate for the loss of linearity, thereby improving convergence properties [48]. In smooth Riemannian optimization, various methods have been developed, such as Riemannian gradient descent [23, 1, 4, 9]; nonlinear conjugate gradients [61]; and variance-reduced stochastic gradients [87, 62]. For further details, please refer to [7] and references therein.
Riemannian bilevel optimization: Research on Riemannian bilevel optimization is limited. [48] investigated hypergradient computation for such problems on Riemannian manifolds and developed deterministic and stochastic algorithms, namely RieBO and RieSBO, for solving Riemannian bilevel and stochastic bilevel optimization. In parallel, [29] proposed a framework called RHGD for Riemannian bilevel optimization, offering multiple hypergradient estimation strategies accompanied by convergence and complexity analyses. The authors also extended their framework to Riemannian minimax and compositional optimization problems. By leveraging value-function reformulation and the Lagrange method, [19] introduced a fully stochastic first-order approach applicable to scenarios where both objectives are stochastic, deterministic, or mixed.
Adaptive bilevel optimization: The closest related approach to our method is the D-TFBO proposed in [81], which tackles Euclidean bilevel optimization by employing a similar adaptive step size strategy. While some aspects of Euclidean analysis extend directly to Riemannian settings, geometric curvature introduces distortions that create unique analytical challenges. We derive convergence guarantees similar to those of D-TFBO and enhance our algorithm’s efficiency by replacing exponential mapping with retraction mapping, thereby reducing computational overhead. To our knowledge, this work presents the first fully adaptive method with theoretical convergence guarantees for solving Riemannian bilevel optimization problems.
Appendix B Additional preliminaries for Section 2
The Lipschitzness of the functions and operators in the Riemannian manifolds are defined as follows:
Definition B.1.
[29, Definition 1] For any $x,x_{1},x_{2}\in{\mathcal{U}}_{x},y,y_{1},y_{2}\in{\mathcal{U}}_{y}$, where ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$,
  1. (i)
    a function $f:{\mathcal{M}}_{x}\rightarrow{\mathbb{R}}$ is said to have $L$-Lipschitz Riemannian gradient in ${\mathcal{U}}_{x}$ if $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}f(x_{1})-{\mathcal{G}}f(x_{2})\|_{x_{2}}\leq Ld(x_{1},x_{2})$.
  2. (ii)
    a bi-function $f:{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}\rightarrow{\mathbb{R}}$ is said to have $L$-Lipschitz Riemannian gradient in ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}$ if $\|{\mathcal{P}}_{y_{1}}^{y_{2}}{\mathcal{G}}_{y}f(x,y_{1})-{\mathcal{G}}_{y}f(x,y_{2})\|_{y_{2}}\leq Ld(y_{1},y_{2})$, $\|{\mathcal{G}}_{x}f(x,y_{1})-{\mathcal{G}}_{x}f(x,y_{2})\|_{x}\leq Ld(y_{1},y_{2})$, $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y)-{\mathcal{G}}_{x}f(x_{2},y)\|_{x_{2}}\leq Ld(x_{1},x_{2})$ and $\|{\mathcal{G}}_{y}f(x_{1},y)-{\mathcal{G}}_{y}f(x_{2},y)\|_{y}\leq Ld(x_{1},x_{2})$.
  3. (iii)
    a linear operator ${\mathcal{G}}(x,y):T_{y}{\mathcal{M}}_{y}\rightarrow T_{x}{\mathcal{M}}_{x}$ (e.g. ${\mathcal{G}}_{xy}g(x,y)$), is said to be $\rho$-Lipschitz in ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}$ if $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}(x_{1},y)-{\mathcal{G}}(x_{2},y)\|_{x_{2}}\leq\rho\,d(x_{1},x_{2})$ and $\|{\mathcal{G}}(x,y_{1})-{\mathcal{G}}(x,y_{2}){\mathcal{P}}_{y_{1}}^{y_{2}}\|_{x}\leq\rho\,d(y_{1},y_{2})$.
  4. (iv)
    a linear operator ${\mathcal{H}}(x,y):T_{y}{\mathcal{M}}_{y}\rightarrow T_{y}{\mathcal{M}}_{y}$ (e.g. ${\mathcal{H}}_{y}g(x,y)$), is said to be $\rho$-Lipschitz in ${\mathcal{U}}_{x}\times{\mathcal{U}}_{y}$ if $\|{\mathcal{P}}_{y_{1}}^{y_{2}}{\mathcal{H}}(x,y_{1}){\mathcal{P}}_{y_{2}}^{y_{1}}-{\mathcal{H}}(x,y_{2})\|_{y_{2}}\leq\rho\,d(y_{1},y_{2})$ and $\|{\mathcal{H}}(x_{1},y)-{\mathcal{H}}(x_{2},y)\|_{y}\leq\rho\,d(x_{1},x_{2})$.
Influenced by the curvature of the Riemannian manifolds, different from the distance defined in the Euclidean space, we have the following trigonometric distance bound over the Riemannian manifolds.
Lemma B.1 (Trigonometric distance bound [86, 87, 26]).
Let $x_{a},x_{b},x_{c}\in{\mathcal{U}}\subseteq{\mathcal{M}}$ and denote $a=d(x_{b},x_{c})$, $b=d(x_{a},x_{c})$ and $c=d(x_{a},x_{b})$ as the geodesic side lengths. Then, it holds that
$a^{2}\leq\zeta(\tau,c)b^{2}+c^{2}-2\left\langle{\rm Exp}_{x_{a}}^{-1}(x_{b}),{\rm Exp}_{x_{a}}^{-1}(x_{c})\right\rangle_{x_{a}}$
where $\zeta(\tau,c)=\frac{\sqrt{|\tau|}c}{\tanh(\sqrt{|\tau|}c)}$ if $\tau<0$ and $\zeta(\tau,c)=1$ if $\tau\geq 0$, and $\tau$ denotes the lower bound of the sectional curvature of ${\mathcal{U}}$ [59, Section 3.1.3].
Appendix C Tangent space conjugate gradient algorithm
In this section, we introduce the tangent-space conjugate gradient method [69, 7], as required in Section 3.1.
Algorithm 2 Tangent Space Conjugate Gradient $v^{n}={\rm TSCG}({\mathcal{H}}_{y}g(x,\hat{y}),{\mathcal{G}}_{y}f(x,\hat{y}),v^{0},\epsilon_{v}$)
1:Let $r^{0}={\mathcal{G}}_{y}f(x,\hat{y})\in T_{y}{\mathcal{M}}_{y},p^{0}=r^{0}$.
2:while ${\mathcal{H}}_{y}g(x,\hat{y})[v^{n}]-{\mathcal{G}}_{y}f(x,\hat{y})>\epsilon_{v}$ do
3:     Compute ${\mathcal{H}}_{y}g(x,\hat{y})[p^{n}]$.
4:     $a^{n+1}=\frac{\|r^{n}\|^{2}_{\hat{y}}}{\langle p^{n},{\mathcal{H}}_{y}g(x,\hat{y})[p^{n}]\rangle_{\hat{y}}}$,
5:     $v^{n+1}=v^{n}+a^{n+1}p^{n}$,
6:     $r^{n+1}=r^{n}-a^{n+1}{\mathcal{H}}_{y}g(x,\hat{y})[p^{n}]$,
7:     $b^{n+1}=\frac{\|r^{n+1}\|_{\hat{y}}^{2}}{\|r^{n}\|_{\hat{y}}^{2}}$,
8:     $p^{n+1}=r^{n+1}+b^{n+1}p^{n}$,
9:     $n=n+1$.
10:end while
Appendix D Additional results of Section 3.4
Algorithm 3 AdaRHD with Retraction (AdaRHD-R)
1:Initia points $x_{0}\in{\mathcal{M}}_{x},y_{0}\in{\mathcal{M}}_{y}$, and $v_{0}\in T_{y_{0}}{\mathcal{M}}$, initial step sizes $a_{0}>0$, $b_{0}>0$, and $c_{0}>0$, total iterations $T$, error tolerances $\epsilon_{y}=\epsilon_{v}=\frac{1}{T}$.
2:for $t=0,1,2,...,T-1$ do
3:     Set $p=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
4:     while $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}>{\epsilon_{y}}$ do
5:         $b_{k+1}^{2}=b_{k}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}$,
6:         $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}(-\frac{1}{b_{k+1}}\,{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,   $p=p+1.$
7:     end while
8:     $K_{t}=k$
9:     Set $n=0$ and $v_{t}^{0}={\mathcal{P}}_{y_{t-1}^{K_{t-1}}}^{y_{t}^{K_{t}}}v_{t-1}^{N_{t-1}}$ if $t>0$ and $v_{0}$ otherwise.
10:     while $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}>{\epsilon_{v}}$ do
11:         $c_{n+1}^{2}=c_{n}^{2}+\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}$,
12:         $v_{t}^{n+1}=v_{t}^{n}-\frac{1}{c_{n+1}}\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})$$\triangleright$ Gradient descent
13:         $n=n+1$.
14:     end while
[…]
16:     $N_{t}=n$,
17:     $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}^{2}_{xy}g(x_{t},y_{t}^{K_{t}})[v_{t}^{N_{t}}]$,
18:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,   $x_{t+1}={\rm Retr}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
19:end for
Since the retraction mapping only affects the total number of iterations required for the lower-level problem in Problem (1) to converge, without influencing the error bounds, the hypergradient approximation error bound (Lemma 3.1) remains consistent with that in Algorithm 1. Moreover, analogous to Proposition 3.2, we establish the following upper bounds for the total iterations required to solve the lower-level problem in Problem (1) and the linear system (4).
Proposition D.1.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, and 3.6 hold. Then, for any $0\leq t\leq T$, the iterations $K_{t}$ and $N_{t}$ required in Algorithm 3 satisfy:
$K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}+\frac{1}{({\mu}/{\bar{b}}-\bar{\zeta}{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{\tilde{b}}{\epsilon_{y}}\right).$
AdaRHD-GD:
$N_{t}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}+\frac{\bar{c}}{\mu}\log\left(\frac{\tilde{c}}{\epsilon_{v}}\right),$
where $C_{b},C_{c}$, $\bar{b}$, $\bar{c}$, $\tilde{b}$, $\tilde{c}$, and $\bar{\zeta}$ are constants defined in Appendix H.
AdaRHD-CG:
$N_{t}\leq\frac{1}{2}{\log\left(\frac{4L_{g}^{3}l_{f}^{2}}{\mu^{3}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{L_{g}/\mu}+1}{\sqrt{L_{g}/\mu}-1}\right)}.$
Appendix E Extend to Riemannian mix-max problems
Riemannian min-max problems [40, 35, 28, 88, 77, 27, 55, 34] have gained increasing interest recently. The problem form of it is
$\min_{x\in{\mathcal{M}}_{x}}\max_{y\in{\mathcal{M}}_{y}}f(x,y),$ (6)
which can be seen as a special case of Riemannian bilevel optimization problem (1) with $g(x,y)=-f(x,y)$.
If $f$ is geodesic strongly convex w.r.t. $y$ over ${\mathcal{M}}_{y}$, the Riemannian hypergradient (3) reduced to ${\mathcal{G}}F(x)={\mathcal{G}}_{x}f(x,y^{*}(x))$, and the approximate Riemannian hypergradient becomes $\widehat{{\mathcal{G}}}F(x,\hat{y})={\mathcal{G}}_{x}f(x,\hat{y})$. Then, the pseudocode of the adaptive method for solving Problem (6) is provided in Algorithm 4.
Algorithm 4 Riemannian bilevel solver for min-max optimization
1:Initia points $x_{0}\in{\mathcal{M}}_{x},y_{0}\in{\mathcal{M}}_{y}$, initial step sizes $a_{0}>0$, $b_{0}>0$, total iterations $T$, error tolerances $\epsilon_{y}=\frac{1}{T}$.
2:for $t=0,1,2,...,T-1$ do
3:     Set $p=0$ and $y_{t}^{0}=y_{t-1}^{K_{t-1}}$ if $t>0$ and $y_{0}$ otherwise.
4:     while $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}>{\epsilon_{y}}$ do
5:         $b_{k+1}^{2}=b_{k}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}$,
6:         $y_{t}^{k+1}={\rm Exp}_{y_{t}^{k}}(-\frac{1}{b_{k+1}},{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$,
7:         Or $y_{t}^{k+1}={\rm Retr}_{y_{t}^{k}}(-\frac{1}{b_{k+1}},{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$$\triangleright$ retraction mapping
8:         $p=p+1.$
9:     end while
10:     $K_{t}=p$,
11:     $\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})={\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})$,
12:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}$,
13:     $x_{t+1}={\rm Exp}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}})$,
14:     Or $x_{t+1}={\rm Retr}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}}).$$\triangleright$ retraction mapping
15:end for
Before providing the convergence result of Algorithm 4, we claim the adopted assumptions required to converge as follows.
Assumption E.1.
  1. (i)
    Assumptions 3.1 and 3.4 hold (Assumptions 3.5 and 3.6 hold when employing retraction);
  2. (ii)
    Assumption 3.2 reduces to $f(x,y)$ is continuously differentiable and $-f(x,y)$ is geodesic strongly convex w.r.t. $y$;
  3. (iii)
    Algorithm 3.3 reduces to $f(x,y)$ is $l_{f}$-Lipschitz continuous. The Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous.
We propose the following convergence result of Algorithm 4 without proof.
Theorem E.1.
Suppose that Assumption E.1 holds. Then, the sequence $\{x_{t}\}_{t=0}^{T}$ generated by Algorithm 4 satisfies
$\frac{1}{T}\sum_{t=0}^{T-1}\|{\mathcal{G}}F(x_{t})\|^{2}_{x_{t}}\leq\frac{C_{mm}}{T}=\mathcal{O}\left(\frac{1}{T}\right),$
where $C_{mm}$ is a constant relevant to the geodetic strong convexity, Lipschitzness, and curvature parameters.
Furthermore, the gradient complexity of Algorithm 4 is $G_{f}=O(1/\epsilon^{2})$.
Appendix F Preliminary results for proofs
Lemma F.1.
[78, Lemma 3.2] For any non-negative $\alpha_{1},...,\alpha_{T}$, and $\alpha_{1}\geq 1$, we have
$\sum_{l=1}^{T}\frac{\alpha_{l}}{\sum_{i=1}^{l}\alpha_{i}}\leq\log\left(\sum_{l=1}^{T}\alpha_{l}\right)+1.$
Lemma F.2.
Suppose that Assumptions 3.1, 3.2, and 3.3 holds. Then, the following statements hold,
  1. (1)
    For $x_{1},x_{2}\in{\mathcal{M}}_{x}$, it holds that
    $d(y^{*}(x_{1}),y^{*}(x_{2}))\leq L_{y}d(x_{1},x_{2}),$
    where $L_{y}:=\frac{L_{g}}{\mu}$ and $y^{*}(x)$ is the optimal solution of the lower-level problem in Problem (1);
  2. (2)
    The Riemannian hypergradient ${\mathcal{G}}F(x)$ satsifies $\|{\mathcal{G}}F(x)\|_{x}\leq l_{f}(1+L_{g}/\mu)$, and for $x_{1},x_{2}\in{\mathcal{M}}_{x}$, it holds that
    $\|{\mathcal{G}}F(x_{1})-{\mathcal{P}}_{x_{2}}^{x_{1}}{\mathcal{G}}F(x_{2})\|_{x_{1}}\leq L_{F}d(x_{1},x_{2}),$
    where $L_{F}:=\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)\left(1+\frac{L_{g}}{\mu}\right)^{2}$;
Proof.
The proofs of Lemma F.2 can be derived from [24, Lemma 2.2], here we provide them for completeness.
(1): For $x_{1},x_{2}\in{\mathcal{M}}_{x}$, there exist two geodesics $c_{1}:[0,1]\rightarrow{\mathcal{M}}_{y}$ and $c_{2}:[0,1]\rightarrow{\mathcal{M}}_{x}$ such that $c_{1}(t)=y^{*}(c_{2}(t))$ for $t\in[0,1]$, and $c_{2}(0)=x_{1},c_{2}(1)=x_{2}$. Then, we have
[…]
where the first inequality follows from $Dy^{*}(x):=-{\mathcal{H}}_{y}^{-1}g(x,y^{*}(x)){\mathcal{G}}^{2}_{yx}g(x,y^{*}(x))$.
(2): By the definition of ${\mathcal{G}}F(x)$, from Assumptions 3.2 and 3.3, it holds that
$\|{\mathcal{G}}F(x)\|_{x}=\|{\mathcal{G}}_{x}f(x,y^{*}(x))-{\mathcal{G}}^{2}_{xy}g(x,y^{*}(x))[{\mathcal{H}}_{y}^{-1}g(x,y^{*}(x))[{\mathcal{G}}_{y}f(x,y^{*}(x))]]\|_{x}$
$=$ $\|{\mathcal{G}}_{x}f(x,y^{*}(x))\|_{x}+\|{\mathcal{G}}^{2}_{xy}g(x,y^{*}(x))\|_{y^{*}(x)}\|{\mathcal{H}}_{y}^{-1}g(x,y^{*}(x))\|_{y^{*}(x)}\|{\mathcal{G}}_{y}f(x,y^{*}(x))\|_{y^{*}(x)}\leq l_{f}+\frac{L_{g}}{\mu}l_{f}.$
Furthermore, we have
$\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}F(x_{1})-{\mathcal{G}}F(x_{2})\|_{x_{2}}$
$\leq$ $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{x}f(x_{2},y^{*}(x_{2}))\|_{x_{2}}$
$+\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{xy}^{2}g(x_{2},y^{*}(x_{2})){\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}\|_{x_{2}}\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))\|_{y^{*}(x_{1})}$
$+\|{\mathcal{G}}^{2}_{xy}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}\|{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}.$ (7)
By Assumptions 3.2 and 3.3, for the first term of (F), we have
$\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{x}f(x_{2},y^{*}(x_{2}))\|_{x_{2}}$
$\leq$ $\|{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{2}))\|_{x_{1}}+\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{x}f(x_{1},y^{*}(x_{2}))-{\mathcal{G}}_{x}f(x_{2},y^{*}(x_{2}))\|_{x_{2}}$
$\leq$ $L_{f}d(y^{*}(x_{1}),y^{*}(x_{2}))+L_{f}d(x_{1},x_{2})=\left(\frac{L_{f}L_{g}}{\mu}+L_{f}\right)d(x_{1},x_{2}).$
We then consider the fourth term of (F), we have
$\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))\|_{y^{*}(x_{2})}+\|{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{2}))-{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\left(\frac{L_{f}L_{g}}{\mu}+L_{f}\right)d(x_{1},x_{2}).$
For the second term of (F), we have
$\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{1}))-{\mathcal{G}}_{xy}^{2}g(x_{2},y^{*}(x_{2})){\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}\|_{x_{2}}$
$\leq$ $\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{2}))\|_{x_{2}}+\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}_{xy}^{2}g(x_{1},y^{*}(x_{2}))-{\mathcal{G}}_{xy}^{2}g(x_{2},y^{*}(x_{2}))\|_{x_{2}}$
$\leq$ $(\frac{\rho L_{g}}{\mu}+\rho)d(x_{1},x_{2}).$
For the third term of (F), we have
$\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\|{\mathcal{H}}^{-1}_{y}g(x_{1},y^{*}(x_{1}))-{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1}))\|_{y^{*}(x_{1})}+\|{\mathcal{P}}_{y^{*}(x_{1})}^{y^{*}(x_{2})}{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{1})){\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}-{\mathcal{H}}^{-1}_{y}g(x_{2},y^{*}(x_{2}))\|_{y^{*}(x_{2})}$
$\leq$ $\left(\frac{\rho}{\mu^{2}}+\frac{\rho L_{g}}{\mu^{3}}\right)d(x_{1},x_{2}),$
where the last inequality follows from the operator norm satisfies
$\|A^{-1}-B^{-1}\|\leq\|B^{-1}(B-A)A^{-1}\|\leq\|A^{-1}\|\|B^{-1}\|\|A-B\|.$ (8)
Combining the above inequalities, we have
$\|{\mathcal{P}}_{x_{1}}^{x_{2}}{\mathcal{G}}F(x_{1})-{\mathcal{G}}F(x_{2})\|_{x_{2}}$
$\leq$ $\left(\frac{L_{f}L_{g}}{\mu}+L_{f}\right)d(x_{1},x_{2})+\frac{l_{f}}{\mu}\left(\frac{\rho L_{g}}{\mu}+\rho\right)d(x_{1},x_{2})$
$+L_{g}l_{f}\left(\frac{\rho}{\mu^{2}}+\frac{\rho L_{g}}{\mu^{3}}\right)d(x_{1},x_{2})+\frac{L_{g}}{\mu}\left(\frac{L_{f}L_{g}}{\mu}+L_{f}\right)d(x_{1},x_{2})$
$=$ $(\frac{L_{f}L_{g}}{\mu}+L_{f}+\frac{l_{f}}{\mu}\left(\frac{\rho L_{g}}{\mu}+\rho\right)+L_{g}l_{f}\left(\frac{\rho}{\mu^{2}}+\frac{\rho L_{g}}{\mu^{3}}\right)+\frac{L_{g}}{\mu}\left(\frac{L_{f}L_{g}}{\mu}+L_{f}\right))d(x_{1},x_{2})$
$=$ $\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)\left(1+\frac{L_{g}}{\mu}\right)^{2}d(x_{1},x_{2}).$
The proof is complete. ∎
Denote $v^{*}(x):=\operatorname*{arg\,min}_{v\in T_{y^{*}(x)}{\mathcal{M}}_{y}}R(x,y^{*}(x),v)$ and $\hat{v}^{*}(x,y):=\operatorname*{arg\,min}_{v\in T_{y}{\mathcal{M}}_{y}}R(x,y,v)$.
Lemma F.3.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, the following statements hold,
  1. (1)
    The objective $R(x,y,v)$ is $\mu$-strongly convex and $L_{g}$-smooth w.r.t. $v$;
  2. (2)
    $v^{*}(x)$ satisfies $\|v^{*}(x)\|_{y^{*}(x)}\leq\frac{l_{f}}{\mu}$, and $\hat{v}^{*}(x,y)$ also satisfies $\|\hat{v}^{*}(x,y)\|_{y}\leq\frac{l_{f}}{\mu}$;
  3. (3)
    For $x,x_{1},x_{2}\in{\mathcal{M}}_{x}$ and $y_{1},y_{2}\in{\mathcal{M}}_{y}$, it holds that
    $\left\|v^{*}(x_{1})-{\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}v^{*}(x_{2})\right\|_{y^{*}(x_{1})}\leq L_{v}d(x_{1},x_{2}),~{}\text{and}~{}\|\hat{v}^{*}(x,y_{1})-{\mathcal{P}}_{y_{2}}^{y_{1}}\hat{v}^{*}(x,y_{2})\|_{y_{1}}\leq\hat{L}_{v}d(y_{1},y_{2}),$
    where $L_{v}:=(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})(1+L_{y})$ and $\hat{L}_{v}:=\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}}$.
Proof.
(1): Since $\nabla_{v}^{2}R(x,y,v)={\mathcal{H}}_{y}g(x,y)$, the strong convexity of $R(x,y,v)$ follows from Assumption 3.2. Then, by Assumption 3.3, for $v_{1},v_{2}\in T_{y}{\mathcal{M}}_{y}$, we have
$\|\nabla_{v}R(x,y,v_{1})-\nabla_{v}R(x,y,v_{2})\|_{y}\leq\|{\mathcal{H}}_{y}g(x,y)\|_{y}\|v_{1}-v_{2}\|_{y}\leq L_{g}\|v_{1}-v_{2}\|_{y}.$
(2): For $\hat{v}^{*}(x,y)$, we have
$\nabla_{v}R\left(x,y,\hat{v}^{*}(x,y)\right)={\mathcal{H}}_{y}g(x,y)[\hat{v}^{*}(x,y)]-{\mathcal{G}}_{y}f(x,y)=0,$
which demonstrates that
$\|\hat{v}^{*}(x,y)\|_{y}=\left\|{\mathcal{H}}_{y}^{-1}g(x,y){\mathcal{G}}_{y}f(x,y)\right\|_{y}\leq\left\|{\mathcal{H}}_{y}^{-1}g(x,y)\right\|_{y}\cdot\|{\mathcal{G}}_{y}f(x,y))\|_{y}\leq\frac{l_{f}}{\mu}.$
Since $v^{*}(x)$ satisfies $v^{*}(x)=\hat{v}^{*}(x,y^{*}(x))$, the desired result follows.
(3): By the definition of $v^{*}(x)$, we have
$\left\|v^{*}(x_{1})-{\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}v^{*}(x_{2})\right\|_{y^{*}(x_{1})}$
$\leq$ $\left\|{\mathcal{H}}_{y}^{-1}g(x_{1},y^{*}(x_{1}))[{\mathcal{G}}_{y}f(x_{1},y^{*}(x_{1}))]-{\mathcal{P}}_{y^{*}(x_{2})}^{y^{*}(x_{1})}{\mathcal{H}}_{y}^{-1}g(x_{2},y^{*}(x_{2}))[{\mathcal{G}}_{y}f(x_{2},y^{*}(x_{2}))]\right\|_{y^{*}(x_{1})}.$
The remaining proof is similar to those of Lemma F.2 (2), thus, we omit it.
By the definition of $\hat{v}^{*}(x,y)$, we have
$\|\hat{v}^{*}(x,y_{1})-{\mathcal{P}}_{y_{2}}^{y_{1}}\hat{v}^{*}(x,y_{2})\|_{y_{1}}=\left\|{\mathcal{H}}_{y}^{-1}g(x,y_{1}){\mathcal{G}}_{y}f(x,y_{1})-{\mathcal{P}}_{y_{2}}^{y_{1}}{\mathcal{H}}_{y}^{-1}g(x,y_{2}){\mathcal{G}}_{y}f(x,y_{2})\right\|_{y_{1}}$
$\leq$ $\left\|{\mathcal{H}}_{y}^{-1}g(x,y_{1})\left({\mathcal{G}}_{y}f(x,y_{1})-{\mathcal{P}}_{y_{2}}^{y_{1}}{\mathcal{G}}_{y}f(x,y_{2})\right)\right\|_{y_{1}}+\left\|\left({\mathcal{H}}_{y}^{-1}g(x,y_{1}){\mathcal{P}}_{y_{2}}^{y_{1}}-{\mathcal{P}}_{y_{2}}^{y_{1}}{\mathcal{H}}_{y}^{-1}g(x,y_{2})\right){\mathcal{G}}_{y}f(x,y_{2})\right\|_{y_{1}}$
$\leq$ $\frac{L_{f}}{\mu}d(y_{1},y_{2})+l_{f}\frac{\rho}{\mu^{2}}d(y_{1},y_{2}),$
where the second inequality follows from (8). ∎
Lemma F.4.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, for any $t\geq 0$, we have
$d(y_{t}^{K_{t}},y^{*}(x_{t}))^{2}\leq\frac{\epsilon_{y}}{\mu^{2}},~{}\text{and}~{}\left\|v_{t}^{N_{t}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq\frac{\epsilon_{v}}{\mu^{2}}.$
Proof.
By the terminated conditions of Algorithm 1 for the lower-level problem and the linear system, we have
$\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}^{2}\leq\epsilon_{y},~{}\text{and}~{}\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{y_{t}^{K_{t}}}^{2}\leq\epsilon_{v}.$
By the strong convexity of $g$ and $R$ (cf. Lemma F.3), we have
$d^{2}(y_{t}^{K_{t}},y^{*}(x_{t}))\leq\frac{1}{\mu^{2}}\left\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}})-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}{\mathcal{G}}_{y}g\left(x_{t},y^{*}(x_{t})\right)\right\|_{y_{t}^{K_{t}}}^{2}\leq\frac{\epsilon_{y}}{\mu^{2}}$
and
$\left\|v_{t}^{N_{t}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq\frac{1}{\mu^{2}}\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})-\nabla_{v}R\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{y_{t}^{K_{t}}}^{2}\leq\frac{\epsilon_{v}}{\mu^{2}},$
where we use the definitions $\|{\mathcal{G}}_{y}g(x_{t},y^{*}(x_{t}))\|^{2}=0$ and $\|\nabla_{v}R\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\|^{2}=0$. ∎
Lemma F.5.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, for any $t\geq 0$, we have
$\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|^{2}\leq C_{f}^{2},$
where $C_{f}:=\left(\frac{2L_{g}^{2}\epsilon_{v}}{\mu^{2}}+\frac{4L_{g}^{2}l_{f}^{2}}{\mu^{2}}+4l_{f}^{2}\right)^{\frac{1}{2}}$.
Proof.
By the definition of $\widehat{\mathcal{G}}F$, we have
$\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$
$\leq$ $2\left\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{x_{t}}^{2}+2\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{x_{t}}^{2}$
$=$ $2\left\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})\left(v_{t}^{N_{t}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{x_{t}}^{2}+2\|{\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}$
$\leq$ $2\left\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})\right\|_{x_{t}}^{2}\|v_{t}^{N_{t}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}^{2}+4\|{\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}+4\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\|_{x_{t}}^{2}$
$\leq$ $\frac{2L_{g}^{2}\epsilon_{v}}{\mu^{2}}+\frac{4L_{g}^{2}l_{f}^{2}}{\mu^{2}}+4l_{f}^{2},$
where the last inequality follows froms Assumptions 3.1, 3.2, 3.3, Lemmas F.3 and F.4. ∎
Appendix G Proofs of Section 3.3
G.1 Proof of Lemma 3.1
Proof.
By the definitions of $v_{t}^{*}(x_{t})$ and $\hat{v}^{*}_{t}(x_{t},y_{t}^{K_{t}})$, we have
$\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})-{\mathcal{G}}F(x_{t})\|_{x_{t}}$
$\leq$ $\|{\mathcal{G}}_{x}f(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}_{x}f(x_{t},y^{*}(x_{t}))\|_{x_{t}}+\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})v_{t}^{N_{t}}-{\mathcal{G}}^{2}_{xy}g(x_{t},y^{*}(x_{t}))v_{t}^{*}(x_{t})\|_{x_{t}}$
$\leq$ $L_{f}d(y_{t}^{K_{t}},y^{*}(x_{t}))+\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})\|_{x_{t}}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}}$
$+\|{\mathcal{G}}_{xy}^{2}g(x_{t},y_{t}^{K_{t}})-{\mathcal{G}}_{xy}^{2}g(x_{t},y^{*}(x_{t})){\mathcal{P}}_{y_{t}^{K_{t}}}^{y^{*}(x_{t})}\|_{x_{t}}\|v_{t}^{*}(x_{t})\|_{y^{*}(x_{t})}$
$\leq$ $L_{f}d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}}+\rho\|v_{t}^{*}(x_{t})\|_{y^{*}(x_{t})}d(y_{t}^{K_{t}},y^{*}(x_{t}))$
$\leq$ $(L_{f}+\frac{\rho l_{f}}{\mu})d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}},$ (9)
where the third and fourth inequalities follow from Assumption 3.3 and Lemma F.3 (2), respectively.
For $\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}}$, we have
$\|v_{t}^{N_{t}}-\mathcal{P}_{y_{t}^{K_{t}}}^{y^{*}(x_{t})}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}}\leq$ $\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+\|\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})-\mathcal{P}_{y_{t}^{K_{t}}}^{y^{*}(x_{t})}\hat{v}_{t}^{*}(x_{t},y^{*}(x_{t}))\|_{y_{t}^{K_{t}}}$
$\leq$ $\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})d(y_{t}^{K_{t}},y^{*}(x_{t})),$
where the last inequality follows from Lemma F.3 (3) and the fact that $v_{t}^{*}(x_{t})=\hat{v}_{t}^{*}(x_{t},y^{*}(x_{t}))$.
Combining it with (G.1) demonstrates that
$\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})-{\mathcal{G}}F(x_{t})\|_{x_{t}}\leq\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{K_{t}}}v_{t}^{*}(x_{t})\|_{y_{t}^{K_{t}}}$
$\leq$ $\left(L_{f}+\frac{\rho l_{f}}{\mu}\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\left(\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}+(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})d(y_{t}^{K_{t}},y^{*}(x_{t}))\right)$
$=$ $\left(L_{f}+\frac{l_{f}\rho}{\mu}+L_{g}\left(\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}}\right)\right)d(y_{t}^{K_{t}},y^{*}(x_{t}))+L_{g}\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{K_{t}})\|_{y_{t}^{K_{t}}}.$
The proof is complete. ∎
G.2 Proof of Proposition 3.2
Inspired by Proposition 1 in [81], we first give a result that concerns the step sizes $a_{t}$, $b_{k}$, and $c_{n}$.
Proposition G.1.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Denote $\{T,K,N\}$ as the iterations of $\{x,y,v\}$. Given any constants $C_{a}\geq a_{0}$, $C_{b}\geq b_{0}$, $C_{c}\geq c_{0}$, then, we have
AdaRHD-GD:
  1. (1)
    either $a_{t}\leq C_{a}$ for any $t\leq T$, or $\exists t_{1}\leq T$ such that $a_{t_{1}}\leq C_{a}$, $a_{t_{1}+1}>C_{a}$;
  2. (2)
    either $b_{k}\leq C_{b}$ for any $k\leq K$, or $\exists k_{1}\leq K$ such that $b_{k_{1}}\leq C_{b}$, $b_{k_{1}+1}>C_{b}$;
  3. (3)
    either $c_{n}\leq C_{c}$ for any $n\leq N$, or $\exists n_{1}\leq N$ such that $c_{n_{1}}\leq C_{c}$, $c_{n_{1}+1}>C_{c}$.
We then consider the case where we use the conjugate gradient method to solve the linear system. Therefore, we do not need to consider the step size $c_{n}$.
AdaRHD-CG:
  1. (1)
    either $a_{t}\leq C_{a}$ for any $t\leq T$, or $\exists t_{1}\leq T$ such that $a_{t_{1}}\leq C_{a}$, $a_{t_{1}+1}>C_{a}$;
  2. (2)
    either $b_{k}\leq C_{b}$ for any $k\leq K$, or $\exists k_{1}\leq K$ such that $b_{k_{1}}\leq C_{b}$, $b_{k_{1}+1}>C_{b}$;
G.2.1 Improvement on objective function for one step update
We first define a threshold for $a_{t}$:
$C_{a}:=\max\{2L_{F},a_{0}\}.$ (10)
Lemma G.1.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, we have
$F(x_{t+1})\leq F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left(1-\frac{L_{F}}{a_{t+1}}\right)\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}}.$ (11)
Furthermore, if $t_{1}$ in Proposition G.1 exists, then for any $t\geq t_{1}$, we have
$F(x_{t+1})\leq F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{4a_{t+1}}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}},$ (12)
where $\epsilon_{y,v}:=\frac{\bar{L}^{2}}{\mu^{2}}(\epsilon_{y}+\epsilon_{v})$ and $\bar{L}:=\max\{\sqrt{2}L_{1},\sqrt{2}L_{g}\}$ with $L_{1}$ is defined in Lemma 3.1.
Proof.
Combining Lemma F.2 (2) and Proposition 2.1, we have
$F(x_{t+1})\leq$ $F(x_{t})+\left\langle{\mathcal{G}}F(x_{t}),{\rm Exp}^{-1}_{x_{t}}(x_{t+1})\right\rangle_{x_{t}}+\frac{L_{F}}{2}d(x_{t+1},x_{t})^{2}$
$=$ $F(x_{t})-\frac{1}{a_{t+1}}\left\langle{\mathcal{G}}F(x_{t}),\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\rangle_{x_{t}}+\frac{L_{F}}{2a_{t+1}^{2}}\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
$=$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x)\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
$+\frac{1}{2a_{t+1}}\left\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}+\frac{L_{F}}{2a_{t+1}^{2}}\left\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
$=$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left(1-\frac{L_{F}}{a_{t+1}}\right)\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$
$+\frac{1}{2a_{t+1}}\left\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}.$ (13)
By Lemma 3.1, it holds that
$\left\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}\leq$ $2L_{1}^{2}d^{2}(y_{t}^{K_{t}},y^{*}(x_{t}))+2L_{g}^{2}\left(\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}}\hat{v}^{*}(x_{t},y^{*}(x_{t}))\|_{y_{t}}^{2}\right)$
$\leq$ $\bar{L}^{2}\left(d^{2}(y_{t}^{K_{t}},y^{*}(x_{t}))+\|v_{t}^{N_{t}}-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}}\hat{v}^{*}(x_{t},y^{*}(x_{t}))\|_{y_{t}}^{2}\right).$
Therefore, by Lemma F.4, we have
$\left\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|^{2}\leq\frac{\bar{L}^{2}}{\mu^{2}}(\epsilon_{y}+\epsilon_{v})=\epsilon_{y,v}.$
This Combine (G.2.1) deduces the desired result of (11).
If $t_{1}$ in Proposition G.1 exists, then for $t\geq t_{1}$, we have $a_{t+1}>C_{a}\geq 2L_{F}$. The desired result of (12) follows from (11). The proof is complete. ∎
Then, similar to [81, Lemma 8], we have the following upper bound for the step size $a_{t}$.
Lemma G.2.
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. If $t_{1}$ in Proposition G.1 does not exist, we have $a_{t}\leq C_{a}$ for all $t\leq T$.
If the $t_{1}$ in Proposition G.1 exists, we have
[…]
where
$F_{0}:=2\left(F(x_{0})-F^{*}\right)+\frac{L_{F}C_{a}^{2}}{a_{0}^{2}}.$ (14)
Proof.
If $t_{1}$ in Proposition G.1 does not exist, then for any $t\leq T$, it holds that $a_{t}\leq C_{a}$.
If $t_{1}$ in Proposition G.1 exists, then for any $t<t_{1}$, it holds that $a_{t+1}\leq C_{a}$. By Lemma G.1, for any $t\geq t_{1}$, it holds that
$F(x_{t+1})\leq F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{4a_{t+1}}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}}.$
Removing the nonnegative term $-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}$, we have
$\frac{\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}}{a_{t+1}}\leq 4\left(F(x_{t})-F(x_{t+1})\right)+\frac{2\epsilon_{y,v}}{a_{t+1}}.$ (15)
Summing (15) from $t_{1}$ to $t$, we have
$\sum_{i=t_{1}}^{t}\frac{\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}}{a_{i+1}}\leq 4\sum_{i={t_{1}}}^{t}\left(F(x_{i})-F(x_{i+1})\right)+\sum_{i={t_{1}}}^{i}\frac{2\epsilon_{y,v}}{a_{i+1}}4\left(F(x_{t_{1}})-F(x_{t+1})\right)+\sum_{i={t_{1}}}^{t}\frac{2\epsilon_{y,v}}{a_{t+1}}.$ (16)
For $F(x_{t_{1}})$, by (11), we have
$F(x_{t_{1}})\leq F(x_{0})+\sum_{i=0}^{t_{1}-1}\frac{L_{F}}{2a_{i+1}^{2}}\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}+\sum_{i=0}^{t_{1}-1}\frac{\epsilon_{y,v}}{2a_{i+1}}.$
This combines with (16), by $F(x_{t+1})\geq F^{*}$, we have
$\sum_{i=t_{1}}^{t}\frac{\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}}{a_{i+1}}\leq 4\left(F(x_{0})-F^{*}\right)+\sum_{i=0}^{t_{1}-1}\frac{2L_{F}}{a_{i+1}^{2}}\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}+\sum_{i=0}^{t}\frac{2\epsilon_{y,v}}{a_{i+1}}$
$\leq$ $4\left(F(x_{0})-F^{*}\right)+\frac{2L_{F}\sum_{i=0}^{t_{1}-1}\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}}{a_{0}^{2}}+\sum_{i=0}^{t}\frac{2\epsilon_{y,v}}{a_{i+1}}$
$\leq$ $4\left(F(x_{0})-F^{*}\right)+\frac{2L_{F}a_{t_{1}}^{2}}{a_{0}^{2}}+\frac{2(t+1)\epsilon_{y,v}}{a_{0}}\leq 4\left(F(x_{0})-F^{*}\right)+\frac{2L_{F}C_{a}^{2}}{a_{0}^{2}}+\frac{2(t+1)\epsilon_{y,v}}{a_{0}}.$
Since $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$, we have
$a_{t+1}=$ $a_{t}+\frac{\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}}{a_{t+1}+a_{t}}\leq a_{t}+\frac{\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}}{a_{t+1}}\leq a_{t_{1}}+\sum_{i=t_{1}}^{t}\frac{\|\widehat{\mathcal{G}}F(x_{i},y_{i}^{K_{i}},v_{i}^{N_{i}})\|_{x_{i}}^{2}}{a_{i+1}}$
$\leq$ $C_{a}+4\left(F(x_{0})-F^{*}\right)+\frac{2L_{F}C_{a}^{2}}{a_{0}^{2}}+\frac{2(t+1)\epsilon_{y,v}}{a_{0}}.$ (17)
The proof is complete. ∎
Similar to (10), we define the thresholds for the step sizes $b_{k}$ and $c_{n}$ as follows:
$C_{b}:=\max\{2\zeta L_{g}\frac{L_{g}}{\mu},b_{0}\},~{}~{}C_{c}:=\max\{L_{g},c_{0}\}.$ (18)
Before proving Proposition 3.2, we need two technical lemmas.
Lemma G.3.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, we have
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})).$
Proof.
By Lemma B.1, we have
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{2}{b_{k+1}}\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\rangle_{y_{t}^{k}}$
$\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{2}{b_{k+1}}(g(x_{t},y^{*}(x_{t}))-g(x_{t},y_{t}^{k})-\frac{\mu}{2}d^{2}(y_{t}^{k},y^{*}(x_{t})))$
$\leq$ $(1+\frac{1}{b_{k+1}^{2}}\zeta L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})),$
where the second inequality follows from the geodesic strong convexity of $g$, and the third inequality follows from
$\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}=\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})-{\mathcal{P}}_{y^{*}(x_{t})}^{y_{t}^{k}}{\mathcal{G}}_{y}g(x_{t},y^{*}(x_{t}))\|^{2}_{y_{t}^{k}}\leq L_{g}^{2}d^{2}(y_{t}^{k},y^{*}(x_{t})).$
The proof is complete. ∎
Lemma G.4.
Suppose that Assumptions 3.1, 3.2, and 3.3 hold. Then, we have
$\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}_{y_{t}^{k}}^{-1}(y^{*}(x_{t}))\rangle_{y_{t}^{k}}\leq-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}.$
Proof.
By Assumption 3.3, given any $y\in{\mathcal{M}}_{y}$, it holds that
$g(x_{t},y^{*}(x_{t}))\leq g(x_{t},y)\leq g(x_{t},y_{t}^{k})+\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}_{y_{t}^{k}}^{-1}(y)\rangle_{y_{t}^{k}}+\frac{L_{g}}{2}d^{2}(y_{t}^{k},y).$ (19)
Taking $y={\rm Exp}_{y_{t}^{k}}(-\frac{1}{L_{g}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}))$ in (19) derives
$g(x_{t},y^{*}(x_{t}))\leq g(x_{t},y_{t}^{k})-\frac{1}{L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+\frac{L_{g}}{2}\|\frac{1}{L_{g}}{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}=g(x_{t},y_{t}^{k})-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}},$
which demonstrates that
$g(x_{t},y^{*}(x_{t}))-g(x_{t},y_{t}^{k})\leq-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}.$ (20)
By the geodesic convex of $g$ w.r.t. $y$, we have
$g(x_{t},y^{*}(x_{t}))-g(x_{t},y_{t}^{k})\geq\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}_{y_{t}^{k}}^{-1}(y^{*}(x_{t}))\rangle_{y_{t}^{k}}.$
This combines (20) implies that
$\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}_{y_{t}^{k}}^{-1}(y^{*}(x_{t}))\rangle_{y_{t}^{k}}\leq g(x_{t},y^{*}(x_{t}))-g(x_{t},y_{t}^{k})\leq-\frac{1}{2L_{g}}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}.$
The proof is complete. ∎
Proof of Proposition 3.2
Proof.
For AdaRHD-GD, denote
$\bar{K}:=\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}+\frac{1}{({\mu}/{\bar{b}}-\zeta{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{\tilde{b}}{\epsilon_{y}}\right),$ (21)
and
$\bar{N}_{gd}:=\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}+\frac{\bar{c}}{\mu}\log\left(\frac{\tilde{c}}{\epsilon_{v}}\right),$ (22)
where $\tilde{b}:=\max\{\frac{L_{g}(\bar{b}-C_{b})}{2},1\}$, $\tilde{c}:=\max\{L_{g}(\bar{c}-C_{c}),1\}$, $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right)$ with $C_{f}=\left(\frac{2L_{g}^{2}\epsilon_{v}}{\mu^{2}}+\frac{4L_{g}^{2}l_{f}^{2}}{\mu^{2}}+4l_{f}^{2}\right)^{\frac{1}{2}}$ is defined in Lemma F.5, and $\bar{c}:=C_{c}+L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right)$.
For AdaRHD-CG, denote
$\bar{N}_{cg}:=\frac{1}{2}{\log\left(\frac{4L_{g}^{3}l_{f}^{2}}{\mu^{3}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{L_{g}/\mu}+1}{\sqrt{L_{g}/\mu}-1}\right)}.$
We first show that $K_{t}\leq\bar{K}$ for all $0\leq t\leq T$.
If $k_{1}$ in Proposition G.1 does not exist, it holds that $b_{K_{t}}\leq C_{b}$. By [80, Lemma 2], we must have $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, since $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{k}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ hold for all $k<K^{t}$, we have
$b_{K_{t}}^{2}=$ $b_{K_{t}-1}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}-1})\|_{y_{t}^{K_{t}-1}}^{2}=b_{K_{t}-1}^{2}\left(1+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}-1})\|_{y_{t}^{K_{t}-1}}^{2}}{b_{K_{t}-1}^{2}}\right)$
$\geq$ $b_{0}^{2}\prod_{k=0}^{K_{t}-1}\left(1+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k}^{2}}\right)\geq b_{0}^{2}\left(1+\frac{\epsilon_{y}}{C_{b}^{2}}\right)^{K_{t}}>b_{0}^{2}\left(\left(1+\frac{\epsilon_{y}}{C_{b}^{2}}\right)^{1/\log(1+\epsilon_{y}/C_{b}^{2})}\right)^{\log(C_{b}^{2}/b_{0}^{2})}=C_{b}^{2},$ (23)
which contradicts $b_{K_{t}}\leq C_{b}$.
If $k_{1}$ in Proposition G.1 exists, then, we have $b_{k_{1}}\leq C_{b}$ and $b_{k_{1}+1}>C_{b}$. We first prove $k_{1}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $k_{1}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, similar to (G.2.1), we have
$b_{k_{1}}^{2}\geq b_{0}^{2}\left(1+\frac{\epsilon_{y}}{C_{b}^{2}}\right)^{k_{1}}>C_{b}^{2},$
which contradicts $b_{k_{1}}\leq C_{b}$.
By Lemma B.1 and the update mode of $y_{t}^{k_{1}}$, we have
$d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))$
$\leq$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta\left\|\frac{{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})}{b_{k_{1}}}\right\|^{2}_{y_{t}^{k_{1}-1}}+2\frac{1}{b_{k_{1}}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1}),{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$\overset{(a)}{\leq}$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta\left\|\frac{{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})}{b_{k_{1}}}\right\|_{y_{t}^{k_{1}-1}}^{2}+2\frac{1}{b_{k_{1}}}\left(g(x_{t},y^{*}(x_{t}))-g(x_{t},y_{t}^{k_{1}-1})-\frac{\mu}{2}d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))\right)$
$\leq$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})\|_{y_{t}^{k_{1}-1}}^{2}}{b_{k_{1}}^{2}}\leq d^{2}(y_{t}^{0},y^{*}(x_{t}))+\zeta\sum_{k=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k+1}^{2}}$
$\overset{(b)}{\leq}$ $d^{2}(y_{t-1}^{K_{t-1}},y^{*}(x_{t}))+\zeta\sum_{k=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}/b_{0}^{2}}{\sum_{i=0}^{k}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{i})\|_{y_{t}^{i}}^{2}/b_{0}^{2}}$
$\overset{(c)}{\leq}$ $2d^{2}(y_{t-1}^{K_{t-1}},y^{*}(x_{t-1}))+2d^{2}(y^{*}(x_{t-1}),y^{*}(x_{t}))+\zeta\log\left(\sum_{k=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{0}^{2}}\right)+\zeta$
$\overset{(d)}{\leq}$ $\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}\|\widehat{\mathcal{G}}F(x_{t-1},y_{t-1}^{K_{t-1}},v_{t-1}^{N_{t-1}})\|_{y_{t-1}^{K_{t-1}}}^{2}}{\mu^{2}a_{t}^{2}}+\zeta\log\left(\sum_{k=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{0}^{2}}\right)+\zeta$
$\overset{(e)}{\leq}$ $\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta,$ (24)
where (a) follows from (2), (b) follows from the warm start of $y_{t}^{0}$, (c) follows from [7, Definition 10.1] and Lemma F.1, (d) uses Lemmas F.2 (1) and F.4, (e) follows from Lemma F.5 and $b_{k_{1}}\leq C_{b}$.
For any $K>k_{1}$, by Lemma G.3, we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq$ $\left(1-\left(\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}\right)\right)d^{2}(y_{t}^{K-1},y^{*}(x_{t}))\leq e^{-\left(\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}\right)(K-k_{1})}d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))$
$\leq$ $e^{-\left(\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}\right)(K-k_{1})}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right),$ (25)
where the second inequality follows from $b_{K}\geq C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$ and $1-x\leq e^{-x}$ for $10<x<$ 1, specifically, we have $C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$ and it holds that $10<\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}<$ 1, and the third inequality follows from (G.2.1).
Similar to (G.2.1), we have
$b_{K}=b_{K-1}+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}+b_{K-1}}\leq b_{k_{1}}+\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k+1}}.$ (26)
Moreover, by the update mode of $y_{t}^{K}$ and Lemma B.1, we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq d^{2}(y_{t}^{K-1},y^{*}(x_{t}))+\zeta\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}^{2}}+\frac{2\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1}),{\rm Exp}^{-1}_{y_{t}^{K-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{K-1}}}{b_{K}}$
$\leq$ $d^{2}(y_{t}^{K-1},y^{*}(x_{t}))+\zeta\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}^{2}}-\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}L_{g}}$
$\leq$ $d^{2}(y_{t}^{K-1},y^{*}(x_{t}))+\zeta\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{2b_{K}\zeta L_{g}\frac{L_{g}}{\mu}}-\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}L_{g}}$
$\leq$ $d^{2}(y_{t}^{K-1},y^{*}(x_{t}))-\frac{(1-{\mu}/{2L_{g}})}{L_{g}}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|^{2}}{b_{K}}\leq d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))-\frac{(1-{\mu}/{2L_{g}})}{L_{g}}\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}}{b_{k+1}},$ (27)
where the second and third inequalities follow from Lemma G.4 and $b_{K}\geq C_{b}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, respectively.
By (G.2.1), we have
$(1-\frac{\mu}{2L_{g}})\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k+1}}\leq$ $L_{g}\left(d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))-d^{2}(y_{t}^{K},y^{*}(x_{t}))\right)\leq L_{g}d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))$
$\leq$ $L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right).$
Then, by (26), $\frac{1}{1-\mu/2L_{g}}\leq 2$, and $b_{k_{1}}\leq C_{b}$, we have
$b_{K}\leq C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right).$
Denote $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right)$. Then, under (G.2.1), considering the monotonicity of $\frac{\mu}{b}-\zeta\frac{L_{g}^{2}}{b^{2}}$ w.r.t. $b$ when $b\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, it holds that $0<\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}\leq\frac{\mu}{b_{K}}-\zeta\frac{L_{g}^{2}}{b_{K}^{2}}$ as $\bar{b}\geq b_{K}\geq 2\zeta L_{g}\frac{L_{g}}{\mu}$, we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){(K-k_{1})}}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\zeta\log\left(\frac{C_{b}}{b_{0}}\right)+\zeta\right)=e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){(K-k_{1})}}\frac{\bar{b}-C_{b}}{2L_{g}}.$ (28)
Since $k_{1}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, by the definition of $\bar{K}$ in (21), it holds that $\bar{K}>k_{1}$ as $\tilde{b}\geq 1>\epsilon_{y}$. If $\frac{L_{g}(\bar{b}-C_{b})}{2}\leq 1$, it holds that $\tilde{b}=1$ in (21). Then, by (28), we have
$\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{\bar{K}})\|_{y_{t}^{\bar{K}}}^{2}\leq$ $L_{g}^{2}d^{2}(y_{t}^{\bar{K}},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){(\bar{K}-k_{1})}}\frac{L_{g}(\bar{b}-C_{b})}{2}$
$\overset{(a)}{\leq}$ $e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){\frac{1}{({\mu}/{\bar{b}}-\zeta{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{1}{\epsilon_{y}}\right)}}\frac{L_{g}(\bar{b}-C_{b})}{2}\overset{(b)}{\leq}e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){\frac{1}{({\mu}/{\bar{b}}-\zeta{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{1}{\epsilon_{y}}\right)}}=\epsilon_{y},$ (29)
where (a) follows from the definition of $\bar{K}$, and (b) follows from $\frac{L_{g}(\bar{b}-C_{b})}{2}\leq 1$.
If $\frac{L_{g}(\bar{b}-C_{b})}{2}>1$, it holds that $\tilde{b}=\frac{L_{g}(\bar{b}-C_{b})}{2}$ in (21), and we have
$\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{\bar{K}})\|_{y_{t}^{\bar{K}}}^{2}\leq$ $L_{g}^{2}d^{2}(y_{t}^{\bar{K}},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){(\bar{K}-k_{1})}}\frac{L_{g}(\bar{b}-C_{b})}{2}$
$\leq$ $e^{-(\frac{\mu}{\bar{b}}-\zeta\frac{L_{g}^{2}}{\bar{b}^{2}}){\frac{1}{({\mu}/{\bar{b}}-\zeta{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{{L_{g}(\bar{b}-C_{b})}/{2}}{\epsilon_{y}}\right)}}\frac{L_{g}(\bar{b}-C_{b})}{2}=\epsilon_{y},$ (30)
Above all, we conclude that after at most $\bar{K}$ iterations, the condition $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{\bar{K}})\|_{y_{t}^{\bar{K}}}^{2}\leq\epsilon_{y}$ is satisfied, i.e., $K_{t}\leq\bar{K}$ holds for all $t\geq 0$. We complete the proof.
AdaRHD-GD: We show that $N_{t}\leq\bar{N}_{gd}$ for all $0\leq t\leq T$.
If $n_{1}$ in Proposition G.1 does not exist, we haves $c_{N_{t}}\leq C_{c}$. Similar to $K_{t}$ of the lower-level problem, we have $N_{t}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$. If $N_{t}>\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$, by $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}\geq\epsilon_{v}$ and $c_{n}\leq C_{c}$ hold for all $n<N_{t}$, we have
$c_{N_{t}}^{2}=$ $c_{N_{t}-1}^{2}+\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}-1})\|_{y_{t}^{K_{t}}}^{2}=c_{N_{t}-1}^{2}\left(1+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}-1})\|_{y_{t}^{K_{t}}}^{2}}{c_{N_{t}-1}^{2}}\right)$
$\geq$ $c_{0}^{2}\prod_{n=0}^{N_{t}-1}\left(1+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}}{c_{n}^{2}}\right)\geq c_{0}^{2}\left(1+\frac{\epsilon_{v}}{C_{c}^{2}}\right)^{N_{t}}>C_{c}^{2},$
which contradicts $c_{N_{t}}\leq C_{c}$.
If $n_{1}$ in Proposition G.1 exists and $N_{t}\geq n_{1}$. Then, it holds that $c_{n_{1}}\leq C_{c}$ and $c_{n_{1}+1}>C_{c}$. Similarly, we have $n_{1}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$.
By the update mode of $v_{t}^{n_{1}}$ and the definition of $\hat{v}^{*}(x_{t},y_{t}^{K_{t}})=\operatorname*{arg\,min}_{v\in T_{y_{t}^{K_{t}}}{\mathcal{M}}_{y}}R(x_{t},y_{t}^{K_{t}},v)$, we have
$\left\|v_{t}^{n_{1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}=\left\|v_{t}^{n_{1}-1}-\frac{\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})}{c_{n_{1}}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}$
$=$ $\left\|v_{t}^{n_{1}-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\left\|\frac{\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})}{c_{n_{1}}}\right\|_{y_{t}^{K_{t}}}^{2}$
$-\frac{2}{c_{n_{1}}}\left\langle v_{t}^{n_{1}-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}}),\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})\right\rangle_{y_{t}^{K_{t}}}$
$\overset{(a)}{\leq}$ $\left\|v_{t}^{n_{1}-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\left\|\frac{\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})}{c_{n_{1}}}\right\|_{y_{t}^{K_{t}}}^{2}$
$-\frac{2}{c_{n_{1}}L_{g}}\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})-\nabla_{v}R\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{y_{t}^{K_{t}}}^{2}$
$\leq$ $\left\|v_{t}^{n_{1}-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\left\|\frac{\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n_{1}-1})}{c_{n_{1}}}\right\|_{y_{t}^{K_{t}}}^{2}$
$\leq$ $\left\|v_{t}^{0}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\sum_{n=0}^{n_{1}-1}\left\|\frac{\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})}{c_{n_{1}}}\right\|_{y_{t}^{K_{t}}}^{2}$
$\overset{(b)}{\leq}$ $\left\|{\mathcal{P}}_{y_{t-1}^{K_{t-1}}}^{y_{t}^{K_{t}}}v_{t-1}^{N_{t-1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\sum_{n=0}^{n_{1}-1}\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}/c_{0}^{2}}{\sum_{i=0}^{n}\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{i})\|_{y_{t}^{K_{t}}}^{2}/c_{0}^{2}}$
$\overset{(c)}{\leq}$ $2\left\|v_{t-1}^{N_{t-1}}-\hat{v}^{*}(x_{t-1},y_{t-1}^{K_{t-1}})\right\|_{y_{t-1}^{K_{t-1}}}^{2}+2\left\|\hat{v}^{*}(x_{t-1},y_{t-1}^{K_{t-1}})-{\mathcal{P}}_{y_{t}^{K_{t}}}^{y_{t-1}^{K_{t-1}}}\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t-1}^{K_{t-1}}}^{2}$
$+\log\left(\sum_{n=0}^{n_{1}-1}\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}/c_{0}^{2}\right)+1$
$\leq$ $2\left\|v_{t-1}^{K_{t-1}}-\hat{v}^{*}(x_{t-1},y_{t-1}^{K_{t-1}})\right\|_{y_{t-1}^{K_{t-1}}}^{2}+4\left\|\hat{v}^{*}(x_{t-1},y_{t-1}^{K_{t-1}})\right\|_{y_{t-1}^{K_{t-1}}}^{2}+4\left\|\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}$
$+\log\left(\sum_{n=0}^{n_{1}-1}\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}/c_{0}^{2}\right)+1$
$\overset{(d)}{\leq}$ $\frac{2\epsilon_{v}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1,$ (31)
where (a) follows from Baillon-Haddad Theorem [3, Theorem 5.8 (iv)], i.e., for $y\in{\mathcal{M}}_{y}$ and $v\in T_{y}{\mathcal{M}}_{y}$, it holds that
$\left\langle v-\hat{v}^{*}(x_{t},y),\nabla_{v}R(x_{t},y,v)-\nabla_{v}R(x_{t},y,\hat{v}^{*}(x_{t},y))\right\rangle_{y}\geq\frac{1}{L_{g}}\left\|\nabla_{v}R(x_{t},y,v)-\nabla_{v}R\left(x_{t},y,\hat{v}^{*}(x_{t},y)\right)\right\|_{y}^{2},$ (32)
(b) follows from the warm start of $v_{t}^{0}$, (c) uses Lemma F.1, and (d) follows from Lemmas F.3 (2) and F.4, and $c_{n_{1}}\leq C_{c}$.
Then, for all $N>n_{1}$, we have
$\left\|v_{t}^{N}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}$
$=$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\|_{y_{t}^{K_{t}}}^{2}}{c_{n}^{2}}-\frac{2\left\langle v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}}),\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\right\rangle_{y_{t}^{K_{t}}}}{c_{n}}$
$\overset{(a)}{\leq}$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}-\frac{1}{c_{n}}\left(2-\frac{L_{g}}{c_{n}}\right)\left\langle v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}}),\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\right\rangle_{y_{t}^{K_{t}}}$
$\overset{(b)}{\leq}$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}-\frac{1}{c_{n}}\left\langle v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}}),\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\right\rangle_{y_{t}^{K_{t}}}$
$\overset{(c)}{\leq}$ $\left(1-\frac{\mu}{c_{n}}\right)\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\overset{(d)}{\leq}e^{-{\mu(N-n_{1})}/{c_{n}}}\left\|v_{t}^{n_{1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}$
$\overset{(e)}{\leq}$ $e^{-{\mu(N-n_{1})}/{c_{n}}}\left(\frac{2\epsilon_{v}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right),$
where (a) follows from (32), (b) follows from $c_{n}>C_{c}\geq L_{g}$, (c) follows from $\nabla_{v}R(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}}))=0$ and the $\mu$-strong convexity of $R$, (d) follows from $c_{n}\geq C_{c}\geq L_{g}\geq\mu$ and $1-x\leq e^{-x}$ for $10<x<$ 1, and (e) follows from (G.2.1).
Similar to (26), we have
$c_{N}=c_{N-1}+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\|_{y_{t}^{K_{t}}}^{2}}{c_{N}+c_{N-1}}\leq c_{n_{1}}+\sum_{n=n_{1}}^{N-1}\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\|_{y_{t}^{K_{t}}}^{2}}{c_{n+1}}.$ (33)
For the second term of (33), we note that
$\left\|v_{t}^{N}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}$
$=$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\|_{y_{t}^{K_{t}}}^{2}}{c_{N}^{2}}-\frac{2\left\langle v_{t}^{N}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}}),\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\right\rangle_{y_{t}^{K_{t}}}}{c_{N}}$
$\leq$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}+\frac{\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\|_{y_{t}^{K_{t}}}^{2}}{c_{N}^{2}}$
$-\frac{2\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})-\nabla_{v}R\left(x_{t},y_{t}^{K_{t}},\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right)\right\|_{y_{t}^{K_{t}}}^{2}}{c_{N}L_{g}}$
$\leq$ $\left\|v_{t}^{N-1}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}-\frac{\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N-1})\right\|_{y_{t}^{K_{t}}}^{2}}{c_{N}L_{g}}$
$\leq$ $\left\|v_{t}^{n_{1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}-\sum_{n=n_{1}}^{N-1}\frac{\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\right\|_{y_{t}^{K_{t}}}^{2}}{c_{n+1}L_{g}},$
where the first inequality follows from (32), and the second inequality follows from $c_{N}\geq C_{c}\geq L_{g}$.
Then, by (G.2.1), it holds that
$\sum_{n=n_{1}}^{N-1}\frac{\left\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{n})\right\|_{y_{t}^{K_{t}}}^{2}}{c_{n+1}}\leq L_{g}\left(\left\|v_{t}^{n_{1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}-\left\|v_{t}^{N}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\right)$
$\leq$ $L_{g}\left\|v_{t}^{n_{1}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right).$
Combining (33) derives
$c_{n}\leq C_{c}+L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right).$
Denote $\bar{c}:=C_{c}+L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right)$. Then, considering the monotonicity of $\frac{\mu}{c}$ w.r.t. $c$, we have
$\left\|v_{t}^{N}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq e^{-{\mu(N_{t}-n_{1})}/{\bar{c}}}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right)=e^{-{\mu(N_{t}-n_{1})}/{\bar{c}}}\frac{\bar{c}-C_{c}}{L_{g}}.$
Since $n_{1}\leq\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}$, by the definition of $\bar{N}_{gd}$ in (22), it holds that $\bar{N}_{gd}>n_{1}$. Then, similar to (G.2.1) and (G.2.1), we have
$\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{\bar{N}_{gd}})\|_{y_{t}^{K_{t}}}^{2}\leq L_{g}^{2}\left\|v_{t}^{\bar{N}_{gd}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq e^{-{\mu(\bar{N}_{gd}-n_{1})}/{\bar{c}}}L_{g}(\bar{c}-C_{c})\leq\epsilon_{v},$
where the last inequality follows from the definition of $\bar{N}_{gd}$. This indicates that, after at most $\bar{N}_{gd}$ iterations, the condition $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{\bar{N}_{gd}})\|_{y_{t}^{K_{t}}}^{2}\leq\epsilon_{v}$ is satisfied, i.e, $N_{t}\leq\bar{N}_{gd}$ holds for all $t\geq 0$.
AdaRHD-CG: We show that $N_{t}\leq\bar{N}_{cg}$ for all $0\leq t\leq T$.
Denote $\kappa_{g}=\frac{L_{g}}{\mu}$. From [7, Equation (6.19)], given $x_{t}\in{\mathcal{M}}_{x}$ and $y_{t}^{N_{t}}\in{\mathcal{M}}_{y}$, we have
$\|v_{t}^{N_{t}}-\hat{v}_{t}^{*}(x_{t},y_{t}^{N_{t}})\|^{2}_{y_{t}^{N_{t}}}\leq 4\kappa_{g}\left(\frac{\sqrt{\kappa_{g}}-1}{\sqrt{\kappa_{g}}+1}\right)^{2N_{t}}\|\hat{v}_{t}^{0}-\hat{v}_{t}^{*}(x_{t},y_{t}^{N_{t}})\|^{2}_{y_{t}^{N_{t}}}\leq 4\kappa_{g}\left(\frac{\sqrt{\kappa_{g}}-1}{\sqrt{\kappa_{g}}+1}\right)^{2N_{t}}\frac{l_{f}^{2}}{\mu^{2}},$
where the last inequality we use the setting of that $\hat{v}_{t}^{0}=0$ and Lemma F.3 (2). Then we have
$\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{y_{t}^{K_{t}}}^{2}\leq L_{g}^{2}\left\|v_{t}^{N_{t}}-\hat{v}^{*}(x_{t},y_{t}^{K_{t}})\right\|_{y_{t}^{K_{t}}}^{2}\leq 4\kappa_{g}L_{g}^{2}\frac{l_{f}^{2}}{\mu^{2}}\left(\frac{\sqrt{\kappa_{g}}-1}{\sqrt{\kappa_{g}}+1}\right)^{2N_{t}},$
which implies that, after at most $\bar{N}_{cg}$ iterations, it holds that $\|\nabla_{v}R(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{y_{t}^{K_{t}}}^{2}\leq\epsilon_{v}$, where
$\bar{N}_{cg}=\frac{1}{2}{\log\left(\frac{4\kappa_{g}L_{g}^{2}l_{f}^{2}}{\mu^{2}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{\kappa_{g}}+1}{\sqrt{\kappa_{g}}-1}\right)}=\frac{1}{2}{\log\left(\frac{4L_{g}^{3}l_{f}^{2}}{\mu^{3}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{L_{g}/\mu}+1}{\sqrt{L_{g}/\mu}-1}\right)}.$
We complete the proof. ∎
G.3 Proof of Theorem 3.1
Proof.
If $t_{1}$ in Proposition G.1 does not exist, we have $a_{T}\leq C_{a}$. Then, by (11) in Lemma G.1, for $t<T$, we have
$\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}\leq 2\left(F(x_{t})-F(x_{t+1})\right)+\frac{L_{F}}{a_{t+1}^{2}}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{a_{t+1}},$
where $\epsilon_{y,v}$ is defined in Lemma G.1. Summing it from $t=0$ to $T-1$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}\leq$ $\frac{2}{T}\left(F(x_{0})-F(x_{T})\right)+\frac{L_{F}}{a_{0}^{2}}\frac{1}{T}\sum_{t=0}^{T-1}\left\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\right\|_{x_{t}}^{2}+\frac{1}{T}\sum_{t=0}^{T-1}\frac{\epsilon_{y,v}}{a_{t+1}}$
$\leq$ $\frac{1}{T}\left(2\left(F(x_{0})-F^{*}\right)+\frac{L_{F}C_{a}^{2}}{a_{0}^{2}}\right)+\frac{\epsilon_{y,v}}{a_{0}}=\frac{F_{0}}{T}+\frac{\epsilon_{y,v}}{a_{0}},$ (34)
where the second inequality follows from $\sum_{t=0}^{T-1}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}\leq a_{T}^{2}\leq C_{a}^{2}$, and $F_{0}$ is defined in (14).
If $t_{1}$ in Proposition G.1 exists, for any $t<t_{1}$, by (11) in Lemma G.1, we have
$\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}\leq 2\left(F(x_{t})-F(x_{t+1})\right)+\frac{L_{F}}{a_{t+1}^{2}}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{a_{t+1}}.$ (35)
For any $t\geq t_{1}$, by (12) in Lemma G.1, we have
$\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}\leq 2\left(F(x_{t})-F(x_{t+1})\right)+\frac{\epsilon_{y,v}}{a_{t+1}}.$ (36)
Summing (35) and (36) from 00 to $T-1$, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}=$ $\frac{1}{T}\left(\sum_{t=0}^{t_{1}-1}\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}+\sum_{t=t_{1}}^{T-1}\frac{\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}}{a_{t+1}}\right)$
$\leq$ $\frac{1}{T}\left(\sum_{t=0}^{T-1}2\left(F(x_{t})-F(x_{t+1})\right)+\frac{L_{F}}{a_{0}^{2}}\sum_{t=0}^{t_{1}-1}\left\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\right\|_{x_{t}}^{2}+\sum_{t=0}^{T-1}\frac{\epsilon_{y,v}}{a_{t+1}}\right)$
$\leq$ $\frac{1}{T}\left(2\left(F(x_{0})-F^{*}\right)+\frac{L_{F}C_{a}^{2}}{a_{0}^{2}}\right)+\frac{\epsilon_{y,v}}{a_{0}}=\frac{F_{0}}{T}+\frac{\epsilon_{y,v}}{a_{0}},$
where the last inequality follows from Assumption 3.4 and $a_{t_{1}}\leq C_{a}$, and $F_{0}$ is defined in Lemma G.2. This result is the same as (G.3).
Then, since $a_{t+1}\leq a_{T}$, by Lemma G.2, we have
$\frac{1}{T}\sum_{t=0}^{T-1}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}\leq\left(\frac{F_{0}}{T}+\frac{\epsilon_{y,v}}{a_{0}}\right)a_{T}\leq\left(\frac{F_{0}}{T}+\frac{\epsilon_{y,v}}{a_{0}}\right)\left(C_{a}+2F_{0}+\frac{2T\epsilon_{y,v}}{a_{0}}\right).$ (37)
Since $\epsilon_{y}=1/T$ and $\epsilon_{v}=1/T$, by the definition of $\epsilon_{y,v}:=\frac{\bar{L}^{2}}{\mu^{2}}(\epsilon_{y}+\epsilon_{v})$ in Lemma G.1, we have $\epsilon_{y,v}=\frac{2\bar{L}^{2}}{T\mu^{2}}$. Then, by the definition of $F_{0}$, (37) is equivalent to
$\frac{1}{T}\sum_{t=0}^{T-1}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}\leq\frac{\left(F_{0}+{2\bar{L}^{2}}/{a_{0}\mu^{2}}\right)\left(C_{a}+2\left(F_{0}+{2\bar{L}^{2}}/{a_{0}\mu^{2}}\right)\right)}{T}=\mathcal{O}\left(\frac{1}{T}\right).$
Let $C:=(F_{0}+\frac{2\bar{L}^{2}}{a_{0}\mu^{2}})(C_{a}+2(F_{0}+\frac{2\bar{L}^{2}}{a_{0}\mu^{2}}))$. The proof is complete. ∎
G.4 Proof of Corollary 3.1
Proof.
By Theorem 3.1, we have
$T=\mathcal{O}(1/\epsilon).$
For the resolution of the lower-level problem, we have
$K_{t}=\mathcal{O}\left(\frac{1}{\log(1+\epsilon)}\right)=\mathcal{O}\left(\frac{1}{\epsilon}\right).$
Similarly, for the resolution of the linear system, we have
AdaRHD-GD:
$N_{t}=\mathcal{O}\left(\frac{1}{\epsilon}\right).$
AdaRHD-CG:
$N_{t}=\mathcal{O}\left(\log\frac{1}{\epsilon}\right).$
Then, it is evident that the gradient complexities of $f$ and $g$ are $G_{f}=\mathcal{O}(1/\epsilon)$ and $G_{g}=\mathcal{O}(1/\epsilon^{2})$, respectively. The complexities of computing the second-order cross derivative and Hessian-vector product of $g$ are $JV_{g}=\mathcal{O}(1/\epsilon)$, $HV_{g}=\mathcal{O}(1/\epsilon^{2})$ for AdaRHD-GD, and $HV_{g}=\tilde{\mathcal{O}}(1/\epsilon)$ for AdaRHD-CG, respectively. ∎
Appendix H Proofs for Section 3.4
This section provides the proofs of the results from Section 3.4. For consistency, we retain the notations introduced in Section 3.3, such as $C_{a}$, $C_{b}$, $C_{c}$, etc.
Firstly, similar to Proposition G.1, we first give a lemma that concerns the step sizes $a_{1}$, $b_{t}$, and $c_{t}$.
Proposition H.1.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.5, and 3.6 hold. Denote $\{T,K,N\}$ as the iterations of $\{x,y,v\}$. Given any constants $C_{a}\geq a_{0}$, $C_{b}\geq b_{0}$, $C_{c}\geq c_{0}$, then, we have
AdaRHD-GD:
  1. (1)
    either $a_{t}\leq C_{a}$ for any $t\leq T$, or $\exists t_{1}\leq T$ such that $a_{t_{1}}\leq C_{a}$, $a_{t_{1}+1}>C_{a}$;
  2. (2)
    either $b_{t}\leq C_{b}$ for any $t\leq K$, or $\exists k_{1}\leq K$ such that $b_{k_{1}}\leq C_{b}$, $b_{k_{1}+1}>C_{b}$;
  3. (3)
    either $c_{t}\leq C_{c}$ for any $t\leq N$, or $\exists n_{1}\leq N$ such that $c_{n_{1}}\leq C_{c}$, $c_{n_{1}+1}>C_{c}$.
We then consider the case where the conjugate gradient solves the linear system. Therefore, we do not need to consider the step size $c_{t}$.
AdaRHD-CG:
  1. (1)
    either $a_{t}\leq C_{a}$ for any $t\leq T$, or $\exists t_{1}\leq T$ such that $a_{t_{1}}\leq C_{a}$, $a_{t_{1}+1}>C_{a}$;
  2. (2)
    either $b_{t}\leq C_{b}$ for any $t\leq K$, or $\exists k_{1}\leq K$ such that $b_{k_{1}}\leq C_{b}$, $b_{k_{1}+1}>C_{b}$;
H.1 Proof of Proposition D.1
H.1.1 Improvement on objective function for one step update
Similar to Lemma G.1, we have the following lemma.
Lemma H.1.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.5, and 3.6 hold. Then, we have
$F(x_{t+1})\leq F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left(1-\frac{\bar{L}_{F}}{a_{t+1}}\right)\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}}.$
Furthermore, if $t_{1}$ in Proposition H.1 exists, then for $t\geq t_{1}$, we have
$F(x_{t+1})\leq F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{4a_{t+1}}\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}},$ (38)
where $\bar{L}_{F}=(L_{F}c_{u}+2c_{R}(l_{f}+l_{f}L_{g}/\mu)$, $\epsilon_{y,v}=\frac{\bar{L}^{2}}{\mu^{2}}(\epsilon_{y}+\epsilon_{v})$, $\bar{L}:=\max\{\sqrt{2}(L_{g}(1+\frac{L_{f}}{\mu}+\frac{l_{f}\rho}{\mu^{2}})+\frac{l_{f}\rho}{\mu}),\sqrt{2}L_{g}\}$, and $c_{u},c_{R}$ are defined in Assumption 3.6.
Proof.
By Lemma F.2, we have
$F(x_{t+1})\leq F(x_{t})+\left\langle{\mathcal{G}}F(x_{t}),{\rm Exp}^{-1}_{x_{t}}(x_{t+1})\right\rangle_{x_{t}}+\frac{L_{F}}{2}d(x_{t+1},x_{t})^{2}$
$=$ $F(x_{t})+\left\langle{\mathcal{G}}F(x_{t}),{\rm Exp}^{-1}_{x_{t}}(x_{t+1})-{\rm Retr}^{-1}_{x_{t}}(x_{t+1})\right\rangle_{x_{t}}+\left\langle{\mathcal{G}}F(x_{t}),{\rm Retr}^{-1}_{x_{t}}(x_{t+1})\right\rangle_{x_{t}}+\frac{L_{F}}{2}d(x_{t+1},x_{t})^{2}$
$\leq$ $F(x_{t})+c_{R}\left(l_{f}+l_{f}\frac{L_{g}}{\mu}\right)\|{\rm Retr}^{-1}_{x_{t}}(x_{t+1})\|_{x_{t}}^{2}-\frac{1}{a_{t+1}}\left\langle{\mathcal{G}}F(x_{t}),\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\rangle_{x_{t}}$
$+\frac{L_{F}c_{u}}{2a_{t+1}^{2}}\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\|_{x_{t}}^{2}$
$\leq$ $F(x_{t})+\frac{c_{R}\left(l_{f}+l_{f}L_{g}/\mu\right)}{a_{t+1}^{2}}\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\|_{x_{t}}^{2}+\frac{L_{F}c_{u}}{2a_{t+1}^{2}}\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\|_{x_{t}}^{2}$
$-\frac{1}{a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{a_{t+1}}\|\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\|_{x_{t}}^{2}+\frac{1}{a_{t+1}}\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\|_{x_{t}}^{2}$
$=$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left(1-\frac{L_{F}c_{u}+2c_{R}\left(l_{f}+l_{f}L_{g}/\mu\right)}{a_{t+1}}\right)\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$
$+\frac{1}{2a_{t+1}}\left\|{\mathcal{G}}F(x_{t})-\widehat{\mathcal{G}}F\left(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}\right)\right\|_{x_{t}}^{2}$
$\leq$ $F(x_{t})-\frac{1}{2a_{t+1}}\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}-\frac{1}{2a_{t+1}}\left(1-\frac{L_{F}c_{u}+2c_{R}\left(l_{f}+l_{f}L_{g}/\mu\right)}{a_{t+1}}\right)\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}+\frac{\epsilon_{y,v}}{2a_{t+1}},$ (39)
where the second inequality uses Lemma F.2 (2) and Assumption 3.6.
If $t_{1}$ in Proposition H.1 exists, then for $t\geq t_{1}$, we have $a_{t+1}>C_{a}\geq 2(L_{F}c_{u}+2c_{R}(l_{f}+l_{f}L_{g}/\mu))$. The desired result of (38) follows from (H.1.1). The proof is complete. ∎
Similar to (10), we define a threshold for parameters $a_{t}$ when giving an upper bound of the step size $a_{t}$.
$C_{a}:=\max\{2\bar{L}_{F},a_{0}\}.$
where $\bar{L}_{F}=(L_{F}c_{u}+2c_{R}(l_{f}+l_{f}L_{g}/\mu)$ is defined in Lemmas H.1.
Similar to Lemma G.2, we have the following upper bound for the step size $a_{t}$.
Lemma H.2.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, and 3.6 hold. If $t_{1}$ in Proposition H.1 does not exist, we have $a_{t}\leq C_{a}$ for any $t\leq T$.
If there exists $t_{1}\leq T$ described in Proposition H.1, we have
[…]
where we define
$F_{0}:=2\left(F(x_{0})-F^{*}\right)+\frac{\bar{L}_{F}C_{a}^{2}}{a_{0}^{2}}.$ (40)
The proof of Lemma H.2 closely parallels that of Lemma G.2, requiring only the substitution of $L_{F}$ with $\bar{L}_{F}$. Here we omit it.
Before proving Proposition D.1, similar to Lemma G.3, we present the following technical lemma when substituting the exponential mapping with the retraction mapping.
Lemma H.3.
Suppose that Assumptions 3.1, 3.2, 3.3, 3.5, and 3.6 hold. Then, we have
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq(1+\frac{1}{b_{k+1}^{2}}\bar{\zeta}L_{g}^{2}-\frac{\mu}{b_{k+1}})d^{2}(y_{t}^{k},y^{*}(x_{t})),$
where $\bar{\zeta}:=\zeta c_{u}+2\bar{D}c_{R}$.
Proof.
From Lemma B.1, by the definition that ${\rm Retr}^{-1}_{y_{t}^{k}}y_{t}^{k+1}=-\frac{1}{b_{k+1}}{\mathcal{G}}g(x_{t},y_{t}^{k})$ we have
$d^{2}(y_{t}^{k+1},y^{*}(x_{t}))\leq d^{2}(y_{t}^{k},y^{*}(x_{t}))+\zeta d^{2}(y_{t}^{k},y_{t}^{k+1})-2\left\langle{\rm Exp}^{-1}_{y_{t}^{k}}y_{t}^{k+1},{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\right\rangle_{y_{t}^{k}}$
$\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta c_{u}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}-2\left\langle{\rm Exp}^{-1}_{y_{t}^{k}}y_{t}^{k+1}-{\rm Retr}^{-1}_{y_{t}^{k}}y_{t}^{k+1},{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\right\rangle_{y_{t}^{k}}$
$+2\frac{1}{b_{k+1}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\right\rangle_{y_{t}^{k}}$
$\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}\zeta c_{u}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+2\frac{1}{b_{k+1}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\right\rangle_{y_{t}^{k}}$
$+2\bar{D}\|{\rm Exp}^{-1}_{y_{t}^{k}}y_{t}^{k+1}-{\rm Retr}^{-1}_{y_{t}^{k}}y_{t}^{k+1}\|_{y_{t}^{k}}$
$\leq$ $d^{2}(y_{t}^{k},y^{*}(x_{t}))+\frac{1}{b_{k+1}^{2}}(\zeta c_{u}+2\bar{D}c_{R})\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}_{y_{t}^{k}}+2\frac{1}{b_{k+1}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\right\rangle_{y_{t}^{k}}$
$\leq$ $\left(1+\frac{1}{b_{k+1}^{2}}(\zeta c_{u}+2\bar{D}c_{R})L_{g}^{2}-{\frac{\mu}{b_{k+1}}}\right)d^{2}(y_{t}^{k},y^{*}(x_{t})),$ (41)
where the second inequality follows from Assumption 3.6 and
$\langle{\rm Retr}^{-1}_{y_{t}^{k}}y_{t}^{k+1},{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\rangle_{y_{t}^{k}}=-\frac{1}{b_{k+1}}\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k}),{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\rangle_{y_{t}^{k}},$
the third inequality follows from Assumption 3.5 by employing [7, Proposition 10.22], i.e.,
$\|{\rm Exp}^{-1}_{y_{t}^{k}}y^{*}(x_{t})\|=d(y_{t}^{k},y^{*}(x_{t}))\leq\bar{D},$
and the fourth inequality follows from Assumption 3.6. ∎
Similar to (18), we define the following thresholds for the step sizes $b_{k}$ and $c_{n}$.
$C_{b}:=\max\{2\bar{\zeta}L_{g}\frac{L_{g}}{\mu},b_{0}\},~{}~{}C_{c}:=\max\{L_{g},c_{0}\},$
where $\bar{\zeta}=\zeta c_{u}+2\bar{D}c_{R}$ are defined in Lemma H.3.
Proof of Proposition D.1
Proof.
For AdaRHD-GD, denote
$\bar{K}:=\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}+\frac{1}{({\mu}/{\bar{b}}-\bar{\zeta}{L_{g}^{2}}/{\bar{b}^{2}})}\log\left(\frac{\tilde{b}}{\epsilon_{y}}\right),$ (42)
and
$\bar{N}_{gd}:=\frac{\log(C_{c}^{2}/c_{0}^{2})}{\log(1+\epsilon_{v}/C_{c}^{2})}+\frac{\bar{c}}{\mu}\log\left(\frac{\tilde{c}}{\epsilon_{v}}\right),$
where $\tilde{b}:=\max\{\frac{L_{g}(\bar{b}-C_{b})}{2},1\}$, $\tilde{c}:=\max\{L_{g}(\bar{c}-C_{c}),1\}$, $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right)$ with $C_{f}=\left(\frac{2L_{g}^{2}\epsilon_{v}}{\mu^{2}}+\frac{4L_{g}^{2}l_{f}^{2}}{\mu^{2}}+4l_{f}^{2}\right)^{\frac{1}{2}}$ is defined in Lemma F.5, and $\bar{c}:=C_{c}+L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{8l_{f}^{2}}{\mu^{2}}+2\log\left(\frac{C_{c}}{c_{0}}\right)+1\right)$.
For AdaRHD-CG, denote
$\bar{N}_{cg}:=\frac{1}{2}{\log\left(\frac{4L_{g}^{3}l_{f}^{2}}{\mu^{3}\epsilon_{v}}\right)}/{\log\left(\frac{\sqrt{L_{g}/\mu}+1}{\sqrt{L_{g}/\mu}-1}\right)}.$
First, we show that $K_{t}\leq\bar{K}$ for all $0\leq t\leq T$.
If $k_{1}$ in Proposition H.1 does not exist, we have $b_{K_{t}}\leq C_{b}$. By [80, Lemma 2], it holds that $K_{t}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $K_{t}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, by $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{x_{t}}^{2}\geq\epsilon_{y}$ and $b_{k}\leq C_{b}$ holds for all $k<K^{t}$, we have
$b_{K_{t}}^{2}=$ $b_{K_{t}-1}^{2}+\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}-1})\|_{y_{t}^{K_{t}-1}}^{2}=b_{K_{t}-1}^{2}\left(1+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K_{t}-1})\|_{y_{t}^{K_{t}-1}}^{2}}{b_{K_{t}-1}^{2}}\right)$
$\geq$ $b_{0}^{2}\prod_{k=0}^{K_{t}-1}\left(1+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k}^{2}}\right)\geq b_{0}^{2}\left(1+\frac{\epsilon_{y}}{C_{b}^{2}}\right)^{K_{t}}>b_{0}^{2}\left((1+\frac{\epsilon_{y}}{C_{b}^{2}})^{1/\log(1+\epsilon_{y}/C_{b}^{2})}\right)^{\log(C_{b}^{2}/b_{0}^{2})}=C_{b}^{2}.$ (43)
This contradicts $b_{K_{t}}\leq C_{b}$.
If $k_{1}$ in Proposition H.1 exists, we have $b_{k_{1}}\leq C_{b}$ and $b_{k_{1}+1}>C_{b}$. We first prove $k_{1}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$. If $k_{1}>\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, similar to (H.1.1), we have
$b_{k_{1}}^{2}\geq b_{0}^{2}\left(1+\frac{\epsilon_{y}}{C_{b}^{2}}\right)^{k_{1}}>C_{b}^{2},$
which contradicts the setting that $b_{k_{1}}\leq C_{b}$.
By Lemma H.3, we have
$d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))\leq d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta d^{2}(y_{t}^{k_{1}-1},y_{t}^{k_{1}})-2\left\langle{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y_{t}^{k_{1}},{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$\overset{(a)}{\leq}$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta c_{u}\left\|\frac{{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})}{b_{k_{1}}}\right\|^{2}_{y_{t}^{k_{1}-1}}-2\left\langle{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y_{t}^{k_{1}}-{\rm Retr}^{-1}_{y_{t}^{k_{1}-1}}y_{t}^{k_{1}},{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$+2\frac{1}{b_{k_{1}}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1}),{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$\overset{(b)}{\leq}$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\zeta c_{u}\left\|\frac{{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})}{b_{k_{1}}}\right\|^{2}_{y_{t}^{k_{1}-1}}+2\frac{1}{b_{k_{1}}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1}),{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$+2\bar{D}\|{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y_{t}^{k_{1}}-{\rm Retr}^{-1}_{y_{t}^{k_{1}-1}}y_{t}^{k_{1}}\|_{y_{t}^{k_{1}-1}}$
$\overset{(c)}{\leq}$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+(\zeta c_{u}+2\bar{D}c_{R})\left\|\frac{{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})}{b_{k_{1}}}\right\|^{2}_{y_{t}^{k_{1}-1}}+2\frac{1}{b_{k_{1}}}\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1}),{\rm Exp}^{-1}_{y_{t}^{k_{1}-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{k_{1}-1}}$
$\overset{(d)}{\leq}$ $d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))+\bar{\zeta}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k_{1}-1})\|_{y_{t}^{k_{1}-1}}^{2}}{b_{k_{1}}^{2}}\leq d^{2}(y_{t}^{0},y^{*}(x_{t}))+\bar{\zeta}\sum_{p=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{p}}^{2}}{b_{k+1}^{2}}$
$\overset{(e)}{\leq}$ $d^{2}(y_{t-1}^{K_{t}-1},y^{*}(x_{t}))+\bar{\zeta}\sum_{p=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{p}}^{2}/b_{0}^{2}}{\sum_{k=0}^{p}\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}/b_{0}^{2}}$
$\overset{(f)}{\leq}$ $2d^{2}(y_{t-1}^{K_{t}-1},y^{*}(x_{t-1}))+2d^{2}(y^{*}(x_{t-1}),y^{*}(x_{t}))+\bar{\zeta}\log\left(\sum_{p=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{p}}^{2}}{b_{0}^{2}}\right)+\bar{\zeta}$
$\overset{(g)}{\leq}$ $\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}\|\widehat{\mathcal{G}}F(x_{t-1},y_{t-1}^{K_{t-1}},v_{t-1}^{N_{t-1}})\|_{y_{t-1}^{K_{t-1}}}^{2}}{\mu^{2}a_{t}^{2}}+\bar{\zeta}\log\left(\sum_{p=0}^{k_{1}-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{p})\|_{y_{t}^{p}}^{2}}{b_{0}^{2}}\right)+\bar{\zeta}$
$\overset{(h)}{\leq}$ $\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta},$ (44)
where (a) follows from the second inequality of(H.1.1), (b) follows from Assumption 3.5, (c) follows from Assumption 3.6, (d) follows from the convexity of $g$, (e) follows from the warm start of $y_{t}^{0}$, (f) follows from Lemma F.1, (g) follows from Lemmas F.2 (1) and F.4, and (h) follows from Lemma F.5 and $b_{k_{1}}\leq C_{b}$.
For all $K>k_{1}$, by Lemma H.3, we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq$ $\left(1-(\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}})\right)d^{2}(y_{t}^{K-1},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}})(K-k_{1})}d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))$
$\leq$ $e^{-(\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}})(K-k_{1})}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right),$ (45)
where the first inequality follows from $b_{k}\geq C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and $1-m\leq e^{-m}$ for $10<m<$ 1, specifically, we have $C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$ and it holds that $10<\frac{\mu}{b_{k}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{k}^{2}}<$ 1, and the second inequality follows from (H.1.1).
Similar to (26), we have
$b_{K}=b_{K-1}+\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}+b_{K-1}}\leq b_{k_{1}}+\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k+1}}.$ (46)
Similar to (G.2.1), we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq d^{2}(y_{t}^{K-1},y^{*}(x_{t}))+\bar{\zeta}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{b_{K}^{2}}+\frac{2\left\langle{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1}),{\rm Exp}^{-1}_{y_{t}^{K-1}}y^{*}(x_{t})\right\rangle_{y_{t}^{K-1}}}{b_{K}}$
$\leq$ $d^{2}(y_{t}^{K-1},y^{*}(x_{t}))+\bar{\zeta}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}^{2}}{2b_{K}\bar{\zeta}L_{g}\frac{L_{g}}{\mu}}-\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|_{y_{t}^{K-1}}}{b_{K}L_{g}}$
$\leq$ $d^{2}(y_{t}^{K-1},y^{*}(x_{t}))-\frac{(1-{\mu}/{2L_{g}})}{L_{g}}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{K-1})\|^{2}}{b_{K}}\leq d^{2}(y_{t}^{k_{1}-1},y^{*}(x_{t}))-\frac{(1-{\mu}/{2L_{g}})}{L_{g}}\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|^{2}}{b_{k+1}},$
where the second inequality follows from Lemma G.4 and $b_{K}\geq C_{b}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$.
By (H.1.1), we have
$(1-\frac{\mu}{2L_{g}})\sum_{k=k_{1}}^{K-1}\frac{\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{k})\|_{y_{t}^{k}}^{2}}{b_{k+1}}\leq$ $L_{g}\left(d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))-d^{2}(y_{t}^{K},y^{*}(x_{t}))\right)\leq L_{g}d^{2}(y_{t}^{k_{1}},y^{*}(x_{t}))$
$\leq$ $L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right).$
Then, by (46) and $\frac{1}{1-\mu/2L_{g}}\leq 2$, we have
$b_{K}\leq C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right).$
Denote $\bar{b}:=C_{b}+2L_{g}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right)$. Then, under (H.1.1), considering the monotonicity of $\frac{\mu}{b}-\bar{\zeta}\frac{L_{g}^{2}}{b^{2}}$ w.r.t. $b$ when $b\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$, it holds that $\frac{\mu}{\bar{b}}-\bar{\zeta}\frac{L_{g}^{2}}{\bar{b}^{2}}\leq\frac{\mu}{b_{K}}-\bar{\zeta}\frac{L_{g}^{2}}{b_{K}^{2}}$ as $\bar{b}\geq b_{K}\geq 2\bar{\zeta}L_{g}\frac{L_{g}}{\mu}$, we have
$d^{2}(y_{t}^{K},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{\bar{b}}-\bar{\zeta}\frac{L_{g}^{2}}{\bar{b}^{2}}){(K-k_{1})}}\left(\frac{2\epsilon_{y}}{\mu^{2}}+\frac{2L_{g}^{2}C_{f}^{2}}{\mu^{2}a_{0}^{2}}+2\bar{\zeta}\log\left(\frac{C_{b}}{b_{0}}\right)+\bar{\zeta}\right)=e^{-(\frac{\mu}{\bar{b}}-\bar{\zeta}\frac{L_{g}^{2}}{\bar{b}^{2}}){(K-k_{1})}}\frac{\bar{b}-C_{b}}{2L_{g}}.$ (47)
Since $k_{1}\leq\frac{\log(C_{b}^{2}/b_{0}^{2})}{\log(1+\epsilon_{y}/C_{b}^{2})}$, by the definition of $\bar{K}$ in (42), it is obvious that $\bar{K}\geq k_{1}$. Therefore, similar to (G.2.1) and (G.2.1), by (47), we have
$\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{\bar{K}})\|_{y_{t}^{\bar{K}}}^{2}\leq L_{g}^{2}d^{2}(y_{t}^{\bar{K}},y^{*}(x_{t}))\leq e^{-(\frac{\mu}{\bar{b}}-\bar{\zeta}\frac{L_{g}^{2}}{\bar{b}^{2}}){(\bar{K}-k_{1})}}\frac{L_{g}(\bar{b}-C_{b})}{2}\leq\epsilon_{y},$
which indicates that after at most $\bar{K}$ iterations, the condition $\|{\mathcal{G}}_{y}g(x_{t},y_{t}^{\bar{K}})\|_{y_{t}^{\bar{K}}}^{2}\leq\epsilon_{y}$ is satisfied, i.e., $K_{t}\leq\bar{K}$ holds for all $t\geq 0$.
Regarding the total number of iterations $N_{t}$ required to solve the linear system, since solving it does not involve retraction mapping, the total iterations match those of Algorithm 1. The proof is complete. ∎
H.2 Proof of Theorem 3.2
Proof.
The proof of Theorem 3.2 follows a similar approach to Theorem 3.1, requiring only the substitution of $L_{F}$ with $\bar{L}_{F}$ and the definition of the constant $C_{\mathrm{retr}}$ as
$C_{\mathrm{retr}}:=\left(F_{0}+\frac{2\bar{L}_{F}^{2}}{a_{0}\mu^{2}}\right)\left(C_{a}+2\left(F_{0}+\frac{2\bar{L}_{F}^{2}}{a_{0}\mu^{2}}\right)\right),$
where $F_{0}$ is defined in (40) of Lemma H.2. Due to this structural similarity, the remaining proof is omitted. ∎
Appendix I Experimental details
This section expands on the experiments discussed in Section 4 and presents supplemental empirical evaluations. Subsection I.1 details the experimental configurations, while Subsection I.2 includes an additional implementation to further demonstrate the robustness of our proposed algorithm through additional empirical validation.
I.1 Experimental Settings
All implementations are executed using the Geoopt framework [43], matching the reference implementation [29] to guarantee equitable comparison conditions. Furthermore, all the experiments are implemented based on Geoopt [43] and are implemented using Python 3.8 on a Linux server with 256GB RAM and 96-core AMD EPYC 7402 2.8GHz CPU.
I.1.1 Simple problem
For $n=100$, the maximum number of outer iterations in RHGD [29] is set to $200$, whereas for $n=1000$, this number is increased to $400$. In Algorithm 3, the number of outer iterations is set to $T=1000$ for AdaRHD-GD and $T=10000$ for AdaRHD-CG. Following [29], we set $\lambda=0.01$ and fix the step sizes in RHGD as $\eta_{x}=\eta_{y}=0.5$. To ensure consistency, the initial step sizes in Algorithm 3 are set to $a_{0}=b_{0}=c_{0}=2$.
[…]
For solving the lower-level problem, our algorithm follows the stopping criteria outlined in Algorithm 1 or terminates when the number of iterations reaches $\min\{50\lfloor t/5\rfloor,500\}$. For RHGD, the maximum number of iterations is set to either $50$ or $20$.
Moreover, the number of outer iterations in our algorithm is set to $T=1000$ for AdaRHD-CG and $T=10000$ for AdaRHD-GD. Additionally, our algorithm terminates if the approximate hypergradient norm falls below $10^{-4}$. For RHGD, the maximum number of outer iterations is set to $200$.
I.1.2 Deep hyper-representation for classification
The experimental configurations in Subsection 4.2 are maintained consistently across all comparative analyses in Subsection 4.1, except that the number of outer iterations is set to be $T=300$ for AdaRHD-CG and $T=500$ for AdaRHD-GD.
I.2 Additional experiments: hyper-representation problem
I.2.1 Shallow hyper-representation for regression
In this subsection, we adopt a shallow regression framework from [29], incorporating a Stiefel manifold constraint [36, 30, 33] to preserve positive-definiteness in learned representations. The SPD representation is transformed into Euclidean space using the matrix logarithm, thereby establishing a bijection between SPD and symmetric matrices, followed by vectorizing the upper-triangular entries via the ${\rm vec}(\cdot)$ operation. We maintain the problem framework established in Subsection 4.2, replacing the cross-entropy loss with the least-squares loss. The revised problem is as follows:
$\begin{array}[]{lcl}&\min\limits_{\mathbf{A}\in{\rm St}(d,r)}&\sum\limits_{i\in{\mathcal{D}}_{\rm val}}\frac{({\rm vec}(\log(\mathbf{A}^{\top}\mathbf{D}_{i}\mathbf{A}))\beta^{*}(\mathbf{A})-y_{i})^{2}}{2|{\mathcal{D}}_{\rm val}|},\\ &{\rm s.t.}&\beta^{*}(\mathbf{A})=\operatorname*{arg\,min}\limits_{\beta\in{\mathbb{R}}^{{r(r+1)/2}}}\sum\limits_{i\in{\mathcal{D}}_{\rm tr}}\frac{({\rm vec}(\log(\mathbf{A}^{\top}\mathbf{D}_{i}\mathbf{A}))\beta-y_{i})^{2}}{2|{\mathcal{D}}_{\rm tr}|}+\frac{\lambda}{2}\|\beta\|^{2},\end{array}$ (48)
where $\lambda>0$ is the regular parameter. Following [29], we set $d=50$, $r=10$, and $\lambda=0.1$.
In Problem (48), the upper-level objective is optimized on the validation set, whereas the lower-level problem is solved on the training set. Following [29], each element $y_{i}$ is defined by $y_{i}={\rm vec}(\log(\mathbf{A}^{\top}\mathbf{D}_{i}\mathbf{A}))\beta+\epsilon_{i}$, where $\mathbf{A}$ and $\mathbf{D}_{i}$ are randomly generated matrices, $\beta$ is a randomly generated vector, and $\epsilon_{i}$ represents a Gaussian noise term. We generate $n$ samples of $\mathbf{D}_{i}$, evenly divided between the validation and training sets. To align with the step size initialization in [29], we initialize parameters $a_{0}$, $b_{0}$, and $c_{0}$ in our algorithm to 20. Moreover, we configure the number of outer iterations as $T=1000$ for AdaRHD-CG and $T=10000$ for AdaRHD-GD. All other settings align with Section 4.1.
Figure 4: Shallow hyper-representation for regression (Left: $n=200$, Right: $n=1000$).
In Figure 4, we illustrate the validation set loss (Upper Objective) versus time for $n=200$ and $n=1000$, respectively. We observe that AdaRHD-CG converges more effectively and rapidly to a smaller objective loss compared to AdaRHD-GD and RHGD. Furthermore, note that although AdaRHD-GD initially converges more slowly than RHGD (potentially due to imprecise hypergradient estimation affecting the upper-level loss reduction), Figure 4 demonstrates that AdaRHD-GD achieves a lower objective loss than RHGD at specific time intervals. These results collectively highlight the efficiency and robustness of our algorithm, as demonstrated in Section 4.1.
I.2.2 Deep hyper-representation for classification
In this subsection, we further evaluate the efficiency and robustness of our proposed algorithm by extending the deep hyper-representation classification framework introduced in Subsection 4.2 to dataset sampling ratios of $12.5\%$ and $25\%$. For each configuration, five independent trials are executed using distinct random seeds. In each trial, a randomly sampled validation set is reserved for the upper-level problem, with an equally sized training partition allocated to the lower-level task. All remaining experimental parameters align with those defined in Subsection 4.2.
Figure 5: Deep hyper-representation for classification (Left: $n=200$, Right: $n=1000$).
Figure 5 depicts the validation accuracy against outer epochs (iterations) with $12.5\%$ and $25\%$ dataset sampling ratios (due to computational constraints, a subset of the full dataset is utilized to ensure manageable training durations). Table 3 summarizes the computational time required to achieve specified validation accuracy thresholds across algorithms. For each random seed, an independent data subset is sampled. Our algorithm exhibits rapid validation accuracy improvements during initial outer iterations, followed by more gradual advancement in later stages. This behavior likely originates from the inner-loop strategy: early iterations necessitate more intensive inner-loop computations to attain high precision, whereas later stages leverage accumulated progress, thereby requiring fewer inner-loop iterations. In contrast, RHGD-20 inner iterations demonstrate rapid initial accuracy gains but experience pronounced degradation subsequently, particularly at the $25\%$ sampling ratio, where its accuracy drops below that of our method. This instability is inherent in RHGD when fewer inner iterations are employed.
Furthermore, while RHGD-20 exhibits faster time-to-$50\%$-accuracy, its elevated variance (evident in Table 3) confirms inherent instability. Notably, our method achieves target accuracy faster than RHGD despite requiring additional outer iterations. This advantage, maintained under identical initial step sizes, demonstrates both the efficacy and stability of our adaptive approach in addressing Riemannian bilevel optimization problems.
Table 3: Time required for each algorithm to first achieve a specified validation accuracy. “Time(s) to $X\%$” is defined as the time elapsed until the validation accuracy first reaches $X\%$. Values are presented as mean $\pm$ standard deviation across five independent trials using fixed random seeds for all algorithms. Statistically significant results are marked in bold. “Data ratio” refers to the proportion of the full dataset.
AdaRHD-CG AdaRHD-GD RHGD-50 RHGD-20
data ratio = $12.5\%$
Time(s) to $50\%$ $878.60\pm$ 90.44 $1258.83\pm 138.54$ $1188.66\pm 148.46$ 530.03 $\pm$ 306.29
Time(s) to $70\%$ 1666.17 $\pm$ 96.74 $1865.13\pm 426.80$ $3070.22\pm 169.30$ $2085.61\pm 236.33$
data ratio = 25%
Time(s) to $50\%$ 2511.13 $\pm$ 233.16 $3408.80\pm 508.51$ $3944.26\pm 589.55$ $2598.87\pm 645.08$
Appendix J Broader impacts and technical limitations
This paper proposes an algorithm for solving Riemannian bilevel optimization problems. No negative societal impacts are anticipated from this research that warrant disclosure.
Notably, this work focuses exclusively on developing an adaptive double-loop algorithm for deterministic Riemannian bilevel optimization (RBO) problems characterized by geodesically strongly convex lower-level objectives, a limitation stemming from the inherent structural complexities of RBO. Future research directions include: (1) designing single-loop adaptive algorithms; (2) extending the framework to stochastic settings via adaptive step sizes using inverse cumulative stochastic hypergradient norms or other strategies [18, 42, 74, 80, 52, 57]; (3) addressing geodesically convex (non-strongly convex) lower-level objectives through regularization [2] or Riemannian Polyak-Łojasiewicz (PŁ) conditions [11]; and (4) investigating alternative adaptive step size strategies to resolve the factor-$1/\epsilon$ gap in gradient ($G_{g}$) and Hessian-vector product ($HV_{g}$) complexities compared to existing Riemannian bilevel methods [29, 48].